### v1

# Scalable Multi-Agent Path Finding using Collision-Aware Dynamic Alert Mask and a Hybrid Execution Strategy

Abstract
Multi-agent pathfinding (MAPF) remains a critical problem in robotics and autonomous systems, where agents must navigate shared spaces efficiently while avoiding conflicts. Traditional centralized algorithms that have global information, such as Conflict-Based Search (CBS), provide high-quality solutions but become computationally expensive in large-scale scenarios due to the combinatorial explosion of conflicts that need resolution. Conversely, distributed approaches that have local information, particularly learning-based methods, offer better scalability by operating with relaxed information availability, yet often at the cost of solution quality. To address these limitations, we propose a hybrid framework that combines decentralized path planning with a lightweight centralized coordinator. Our framework leverages reinforcement learning (RL) for decentralized planning, enabling agents to adapt their planning based on minimal, targeted alerts–such as static conflict-cell flags or brief conflict tracks–that are dynamically shared information from the central coordinator for effective conflict resolution. We empirically study the effect of the information available to an agent on its planning performance. Our approach reduces the inter-agent information sharing compared to fully centralized and distributed methods, while still consistently finding feasible, collision-free solutions–even in large-scale scenarios having higher agent counts.
1 Introduction
Multi-Agent Path Finding (MAPF) addresses the fundamental challenge of computing collision-free trajectories for multiple agents navigating a shared environment. Its effective solving is crucial for deploying a wide array of multi-agent systems, from automated warehouses and robotic swarms to autonomous vehicle coordination (Sharon et al. 2015). Despite its practical significance, MAPF is computationally demanding, classified as NP-hard in its general form, which can render traditional centralized search methods intractable as the number of agents or the complexity of the environment increases (Ren et al. 2025; Sartoretti et al. 2019).
Figure 1: An example MAPF problem and our four-stage planning pipeline. Left: Three agents (A1, A2, A3) navigate a grid with static obstacles (dark gray). The diagram illustrates a future vertex collision (red X), where two agents would occupy the same cell, and an edge collision (red arrows), where agents would swap adjacent cells. Top Right: The varying levels of information available to Agent A1 under centralized (all agent positions and goals), distributed (nearby agent positions and goals), and decentralized (only nearby agent positions) paradigms. Bottom Right: The four stages of our framework, from initial Path Planning (S1) to Collision Detection (S2), Resolution (S3), and Replanning (S4).
Extensive research in MAPF has explored various coordination paradigms, each with different implications for information availability and system performance. Centralized approaches, such as Conflict-Based Search (CBS) (Sharon et al. 2015) and its efficiency-focused variant ICBS (Boyarski et al. 2015), typically assume full observability of the environment and agent states, helping them generate optimal plans. However, as the number of agents or the size of the environment increases, joint planning and conflict resolution become computationally expensive. Additionally, requiring agents to share full information raises privacy concerns, limiting applicability in confidential settings motivating strategies that reduce reliance on global information. Early decoupled methods like $M^{*}$ (Wagner and Choset 2011) begin with individual plans and only coordinate agents in conflict, offering scalability but often sacrificing optimality or completeness. More recently, Multi-Agent Reinforcement Learning (MARL) has emerged to address coordination and partial observability in dynamic settings. For example, PRIMAL (Sartoretti et al. 2019) trains agents to plan using partial views, learning implicit coordination, while methods like FOLLOWER (Skrynnik et al. 2024) reduce explicit communication by relying on global heuristic maps. These decentralized methods improve scalability but may reduce solution quality and leave conflicts unresolved due to limited information.
To address these limitations, we propose a novel hybrid MAPF framework that combines decentralized planning with a lightweight centralized coordinator, focusing on minimizing inter-agent information sharing while maintaining solution feasibility. Our findings show that minimal, targeted alerts are sufficient, reducing the total information load by an estimated $\sim$93% compared to a continuous-observation distributed paradigm (see Appendix A.1). In our approach, agents primarily rely on decentralized, reinforcement learning-based neural network planners that operate on local observations (e.g., position coordinates), eliminating the need for global information as well as local egocentric maps. A central coordinator oversees agent trajectories, intervening selectively by dynamically sharing targeted information to prompt localized re-planning when conflicts are anticipated. This work’s contributions are centered on this selective coordination strategy. We introduce (1) a novel hybrid framework with this on-demand alert mechanism and provide (2) a comprehensive empirical evaluation demonstrating that our planning policy generalizes robustly from a simple training regimen to larger, more complex environments. Our evaluation is guided by two key research questions: (RQ1) Given the observation constraints of a decentralized setup, can an effective MAPF algorithm be created with one agent knowing nothing about other agents? If not, what information must it need at a minimum? and (RQ2) How does the proposed hybrid method compare to leading alternative search- and learning-based approaches in terms of performance, solution quality, and scalability?
2 Background and Literature Review
2.1 Multi‑Agent Path Finding
Let $G=(V,E)$ be an undirected graph, where $V$ is the set of vertices (grid cells) and $E\subseteq V\times V$ is the set of edges connecting adjacent cells. A team of $n$ agents $A=\{a_{1},\dots,a_{n}\}$ must move from start vertices $s_{i}\in V$ to goal vertices $g_{i}\in V$, where $(s_{i},g_{i})\neq(s_{j},g_{j}),\forall i\neq j:i,j\in\{1,...,n\}$. Time is discretized into steps $t=0,1,2,\dots$, and at each step, an agent may either move along an edge or wait. A path for agent $a_{i}$ is a sequence $\pi_{i}=(v^{i}_{0},v^{i}_{1},\dots,v^{i}_{T_{i}})$ with $v^{i}_{0}=s_{i}$ and $v^{i}_{T_{i}}=g_{i}$. A solution to MAPF is $\Pi=\{\pi_{1},\dots,\pi_{n}\}$, and it is collision‑free if for all $i\neq j$ and all $t$, $v^{i}_{t}\neq v^{j}_{t}$ (vertex collision free) and $(v^{i}_{t},v^{i}_{t+1})\neq(v^{j}_{t+1},v^{j}_{t})$ (edge collision free). The primary goal in standard MAPF is typically to find a path set $\Pi$ that is collision-free (Sharon et al. 2015). Common efficiency objectives include minimizing the makespan, $\max_{i}T_{i}$, minimizing the sum of individual completion times (sum-of-costs), $\sum_{i}T_{i}$, or minimizing the number of collisions.
An important consideration in defining a MAPF problem instance arises from agents potentially reaching their targets at different time steps, thus requiring a clear definition of how an agent behaves after it has reached its target $g_{i}$ at time $T_{i}$, but while other agents may still be en route. Two common settings are used to address this scenario (Stern et al. 2019a). The first setting, typically referred to as stay at target, considers that an agent, upon reaching its goal, remains at that vertex until all other agents in the team have also reached their respective targets. This waiting agent continues to occupy its target vertex, potentially causing conflicts if another agent’s plan requires traversing or occupying that vertex at any time $t\geq T_{i}$ up to the completion time of the entire solution. The second setting, disappear at target, assumes an agent is effectively removed from the environment immediately upon reaching its target. Consequently, its path concludes at time $T_{i}$, and it poses no further collision risk nor occupies any vertex for $t>T_{i}$. In our work, we adopt the second setting, disappear at target. This setting is particularly relevant for applications like drone delivery, where an agent’s task is considered complete upon arrival at a target location.
2.2 Coordination Paradigms
As formalized by Sharon et al. (Sharon et al. 2015), solution approaches to MAPF problems can be categorized based on their coordination strategy. We extend this by focusing on three key elements: state observability (global vs. local), communication (allowed vs. none), and control (centralized vs. decentralized). In centralized paradigms, a global planner with full observability of the entire state controls all agents, and communication is implicit through this central controller. By contrast, other approaches grant agents local control over their actions. Within this category, we draw a key distinction based on communication: In distributed approaches, agents that plan their own paths are allowed to explicitly exchange information, such as local observations, intended paths, or goals, with other agents to achieve cooperative behavior. In decentralized MAPF, as defined in this work, inter-agent communication during execution is eliminated. Agents plan and act entirely independently, relying only on their own local information. Finally, hybrid frameworks, like the one we propose, combine these elements. They typically employ a central coordination mechanism that has full observability to detect conflicts and selectively intervene, while agents otherwise plan in a decentralized manner. In our work, we adopt a hybrid framework with a customized information sharing mechanism as described in Section 3.
2.3 Literature Review
Search-based Methods: Research on MAPF has evolved from centralized to more scalable methods. Early advancements include $M^{*}$, which dynamically couples agent searches upon conflict for optimality and completeness (Wagner and Choset 2011), and Increasing Cost Tree Search (ICTS), which employs cost allocation and then checks for conflict-free solutions from sets of single-agent paths, yielding significant speedups over joint-space search (Sharon et al. 2013). A pivotal development was CBS, where agents independently plan paths using $A^{*}$, and a centralized detector resolves conflicts by branching with new constraints for involved agents, ensuring optimality (Sharon et al. 2015). To enhance scalability while preserving optimality, variants such as ICBS integrate cardinality heuristics and meta-agent merging to accelerate convergence (Boyarski et al. 2015). For very large instances, suboptimal repair strategies like Large Neighborhood Search (LNS) generate initial solutions (often via prioritized planning) and iteratively repair colliding agent subsets under a centralized control loop (Li et al. 2022). However, these methods face scalability challenges as the number of agents increases and often depend on extensive global information for coordination.
Learning-based Methods: In parallel, learning‐based approaches address partial observability and communication constraints. PRIMAL and PRIMAL2 train decentralized policies via a combination of imitation learning from an expert centralized planner and reinforcement learning (PRIMAL2 enhancing local observations for better coordination), effectively implementing a centralized training and decentralized execution (CTDE) paradigm (Sartoretti et al. 2019; Damani et al. 2021b). SCRIMP introduces a transformer‐based communication module enabling agents with limited fields of view to coordinate effectively, achieving decentralized coordination without a central execution-time coordinator (Wang et al. 2023). To function in real‐world settings, each agent must carry onboard sensing, e.g., RGB‐D cameras or LiDAR scanners, and implement an inter-agent communication protocol for goal disclosure, which introduces additional computational overhead and raises privacy concerns. While offering adaptability and decentralized execution, learning-based methods can require significant training data, may struggle with generalization, and often provide weaker guarantees on solution quality or completeness when relying purely on local information.
Hybrid Methods: Hybrid planning‐learning paradigms seek the strengths of both worlds. Authors in (Skrynnik et al. 2024) introduce FOLLOWER, which involves each agent using a congestion‐aware $A^{*}$ planner to determine its subgoals, then executing a fully decentralized RL policy for lifelong replanning. More recently, hybrid frameworks like LNS2+RL, introduced by Li et al., have emerged, augmenting LNS by employing MARL for early, localized conflict resolution, then relying on efficient search-based planning for final refinement (Wang et al. 2025). These existing hybrid approaches, while innovative, can involve complex integrations and may still face trade-offs between optimality, scalability, or specific information dependencies (like global congestion maps (Skrynnik et al. 2024)).
To enable a structured comparison of diverse MAPF techniques (search-based, learning-based, and hybrid), we adopt the four‑stage MAPF pipeline: S1 (Agent Planning), S2 (Collision Detection), S3 (Collision Avoidance Policy), and S4 (Agent Replanning). Within this structure, CBS uses decentralized A* search for S1/S4 and centralized high level solver for collision resolution in S2/S3 for optimality (Sharon et al. 2015). LNS employs a centralized loop for all four stages, managing path generation, detection, repair, and subset replanning, thereby trading optimality for scalability (Li et al. 2022). Learning-based planners like PRIMAL and SCRIMP typically use local observations for decentralized S1-S4 execution, coordinating without a central controller (Sartoretti et al. 2019; Wang et al. 2023); however, practical deployments can require sensors (eg., camera, LiDAR etc.) and communication, adding computational overhead and increase sensing cost. Hybrid methods such as FOLLOWER use global congestion maps with $A^{*}$ search for S1, then decentralized RL policy for S2-S4, often without further communication (Skrynnik et al. 2024). LNS2+RL applies centralized LNS for S1-S3, while S4 uses a hybrid policy (MARL or prioritized planning) for subsets, balancing key trade-offs (Wang et al. 2025). Refer to Table 2 in the Appendix Section A.2 for a detailed categorization of these methodologies.
Together, these works highlight key limitations of existing MAPF paradigms: centralized methods face scalability and privacy concerns; purely learning-based approaches compromise solution quality; and current hybrid solutions often still depend on significant centralized communication or coordination overhead with higher sensing cost. The hybrid framework introduced in Section 3, which is the focus of our work, is specifically designed to mitigate these limitations. It strategically reduces inter-agent information sharing to maintain solution feasibility, thereby addressing the aforementioned information management challenges while also aiming to preclude the additional sensing costs and computational overhead common in learning-based methods.
3 Methodology
3.1 Approach Summary
We propose a hybrid coordination framework that leverages decentralized RL-based path planning together with centralized collision detection and control. While conceptually similar to the high-level loop of Conflict-Based Search (CBS), our framework differs significantly in its resolution step: it does not build a constraint tree or perform backtracking, instead issuing targeted alerts that prompt heuristic-based replanning from the local RL agent. The proposed method’s four-stage MAPF pipeline is detailed below.
S1: Decentralized Path Planning
Each agent $a_{k}$ generates an independent trajectory
$\rho_{k}=\pi_{\theta}(s_{k},g_{k})=(v^{k}_{0},\dots,v^{k}_{\tau_{k}}),$ (1)
where $v^{k}_{0}=s_{k}$, $v^{k}_{\tau_{k}}=g_{k}$, and $\pi_{\theta}$ is a parameterized RL policy trained to minimize path length and local collision risk over a planning horizon $H$. No information about other agents is exchanged during this stage. Let $\tau_{i}$ denote the set of makespan of each agent and $\tau=\{\tau_{1},\dots,\tau_{n}\}$ be the set of all such makespans.
S2–S3: Centralized Collision Detection and Control
A central module takes as input the independent trajectories of all the agents $\rho=\{\rho_{1},\dots,\rho_{n}\}$ along with the makespan set $\tau$, and identifies all vertex and edge conflicts, using
$C(\rho,\tau)=\{(t_{j},\Delta_{c},A_{c})\mid{}(v^{k}_{t_{j}}=v^{l}_{t_{j}}:=\Delta_{c})\;\lor\;$
$((v^{k}_{t_{j}},v^{k}_{t_{j+1}})=(v^{l}_{t_{j}+1},v^{l}_{t_{j}}):=\Delta_{c})\}$ (2)
$A_{c}=\{\{a_{k},\,a_{l}\}:\rho_{k}|\{t_{j}\}=\rho_{l}|\{t_{j}\}=\Delta_{c}\lor\rho_{k}|$
$\{(t_{j},t_{j+1})\}=\rho_{l}|\{(t_{j},t_{j+1})\}=\Delta_{c}$ $,\forall k\neq l\},$ (3)
where $\rho_{k}|\{t_{j}\}$ and $\rho_{k}|\{(t_{j},t_{j+1})\}$ denotes the position of the $k^{th}$ agent at step $t_{j}$ and steps ($t_{j}$, $t_{j+1}$), respectively.
For each conflict $c=(t_{j},\Delta_{c},A_{c})\in C(\rho,\tau)$, which occurs at timestep $t_{j}$, the controller issues an alert $\mathcal{A}$ defined as
$\mathcal{A}(c)=\mu_{c_{k}}=(a_{c_{k}},t_{j-r},\Delta_{c}),r\geq j,$ (4)
where policy $\mu_{c_{k}}$ is used to select an agent $a_{c_{k}}\in A_{c}$. Example policies for this agent selection is described in Section 3.2. Once an alert is issued to an agent, the selected agent is prompted to perform a constrained replanning of its trajectory for a rollout window $\{t^{c_{k}}_{j-r},\dots,t^{c_{k}}_{j},\dots,\tau_{c_{k}}\}$, where $r$ is the rewind window by avoiding the collision set $\Delta_{c}$.
S4: Decentralized Replanning
Upon receiving a collision alert, agent $a_{c_{k}}$ truncates its path at $v^{c_{k}}_{t_{j-r}}$, denoted as
$\rho_{c_{k}|t_{j-r-1}}=(v^{c_{k}}_{0},\dots,v^{c_{k}}_{t_{j-r-1}}).$ (5)
The agent then generates a new path segment $\rho_{c_{k}}^{\prime}$ from $v^{c_{k}}_{t_{j-r}}$ to its goal $g_{c_{k}}$ by applying its RL policy $\pi_{\theta}$. This replanning incorporates new constraints derived from the alert. For static obstacle avoidance (S4.1), the policy is typically derived by incorporating a fixed constraint, wherein the agent is supposed to avoid forbidden cells, as captured in Eq. 6, where $\Delta_{c}$ represents the set of static cells identified in the alert, and a segment of the new path is generated as:
$\rho_{c_{k}}^{\prime}=\pi_{\theta}(v^{c_{k}}_{t_{j-r}},g_{c_{k}}\mid v^{c_{k}}_{t_{i}}\notin\Delta_{c},i\geq j-r).$ (6)
Alternatively, for dynamic obstacle avoidance (S4.2), the replanning must account for the predicted movements of other agents involved in the collision. In this case, the RL policy is conditioned on the sub-paths of the conflicting agents
$\rho_{c_{k}}^{\prime}=\pi_{\theta}(v^{c_{k}}_{t_{j-r}},g_{c_{k}}\mid v^{c_{k}}_{t}\neq v^{c_{l}}_{t},v^{c_{l}}_{t}\in\rho_{c_{l}},$
$\forall l\neq k,a_{c_{l}}\in A_{c},\forall t\in[t_{j-r},t_{j+r}]).$ (7)
The information guiding the replanning–whether it constitutes minimal details such as static obstacle constraints (Eq. 6), or more detailed sub-path information about colliding agents’ paths treated as dynamic obstacles (Eq. 7)–is integrated into the RL agent’s decision-making process (by modifying its state representation, more details in the Section 3.2) to guide it towards a conflict-free solution. The agent’s new complete trajectory $\rho^{new}_{c_{k}}$ is formed by concatenating the initial segment with the newly planned one, given by
$\rho^{new}_{c_{k}}=\rho_{c_{k}|t_{j-r}}\|\rho_{c_{k}}^{\prime}.$ (8)
This updated trajectory is submitted to the central controller. This iterative cycle of detection (S2), control (S3), and selective replanning (S4) continues until a globally conflict-free solution is achieved ($C(\rho,\tau)=\emptyset$).
Our hybrid MAPF framework manages a precise flow of information, essential to support our claims of reduced information sharing and adaptive planning. The process begins with fully decentralized agent planning (S1), where each RL-driven agent uses only its local information before submitting its intended path $\rho_{k}$ to a central coordinator. This coordinator, leveraging its global observation of all submitted paths, performs collision detection (S2) to identify the set of all potential conflicts $C(\rho,\tau)$. Following detection, the central control module (S3) issues targeted alerts; it selects a specific agent $a_{c_{k}}$ and determines the rewind point $t_{j-r}$ from which the agent must replan, effectively issuing an alert $\mathcal{A}(c)$. S3 also governs the level of information to be shared for the subsequent decentralized replanning (S4). Initially, the alerted agent $a_{c_{k}}$ attempts to resolve the conflict using minimal information, typically by treating its own conflicting path segment $\Delta_{c}$ as a static obstacle (S4.1). If this localized, low-information approach proves insufficient, S3 then shares additional information with the agent–specifically, a subset of the path of other directly conflicting agents–for a more informed replan (S4.2). This tiered strategy—from purely local information in S1, to targeted alerts from S3, and then to progressively detailed yet still localized conflict information for S4—ensures that inter-agent information sharing is demand-driven and sparse. Agents communicate their revised plans $\rho^{new}_{c_{k}}$ back to the coordinator, and this iterative cycle of detection and selective, adaptive replanning continues until a globally conflict-free solution is achieved ($C(\rho,\tau)=\emptyset$). This methodology substantiates our framework’s ability to ensure feasibility while operating with significantly reduced information exchange.
3.2 Policy Representation
(S1 & S4) Decentralized Path Planning:
In our proposed framework, we address the challenges of decentralized multi-agent pathfinding in dynamic environments by integrating collision awareness directly into the agent’s observation. Our approach leverages an RL-based decentralized planner that receives a multi-channel grid observation.
Observation Space: Each agent’s observation is a tensor $\mathbf{s}\in\mathbb{R}^{H\times W\times 4}$, where $H$ and $W$ are the height and width of the grid. The four channels correspond to: (a) ObstacleMap - A binary map (1 if a static obstacle is present, 0 otherwise). (b) AgentMap - A binary map marking the agent’s current position. (c) GoalMap - A binary map indicating the goal position. (d) AlertMask - A dynamic collision alert map (initially all zeros, updated during execution). This alert mask is updated online by a centralized collision detection module that simulates timely collision alerts. In addition, the observation includes low-dimensional features comprising a unit vector pointing from the agent’s current position to the goal and the Euclidean distance.
Action Space: The agent operates in a discrete action space, $\mathcal{A}=\{0,1,2,3,4\}$, corresponding to movements in the four cardinal directions (e.g., Up, Down, Left, Right, WAIT).
Reward Structure Our reward function, similar to PRIMAL (Sartoretti et al. 2019), promotes efficient, collision-free navigation and discourages unproductive behaviors. Agents receive +20 for reaching the goal. Penalties include: -3 for collisions with obstacles, -2 for timeout, -0.02 per time step taken (to encourage efficiency), and -0.1 for a ‘WAIT’ action, a penalty that encourages path progression over stationary waiting. An additional -0.05 penalty is applied if the agent is near a dynamic obstacle.
(S2) Rule-Based Collision Detection:
The collision detection module (S2) functions as a deterministic, rule-based system. It takes the set of all current agent trajectories $\rho=\{\rho_{1},\dots,\rho_{n}\}$ and their makespans $\tau$ as input. For each timestep $t=\{0,1,...,T_{M}\}$ upto the maximum makespan ($T_{M}$), the module systematically scans for two primary types of conflicts: vertex conflicts (where $v^{k}_{t}=v^{\ell}_{t}$ for $k\neq\ell$) and edge conflicts (where $(v^{k}_{t},v^{k}_{t+1})=(v^{\ell}_{t+1},v^{\ell}_{t})$ for $k\neq\ell$). Upon detecting a conflict, characterized as $c=(t,v,A_{c})$ where $A_{c}$ is the set of agents involved, S2 reports this conflict $c$ to the S3 control module for resolution. The detection process for all conflicts is computationally efficient, with a time complexity of $O\bigl(\sum_{k}|\rho_{k}|\bigr)$ per cycle, linear in the sum of all agent path lengths.
(S3) Heuristic-Based Collision Avoidance Control:
[…]
4 Experimental Setup
In this section we first describe how we train the per‐agent navigation policy used in stage S1, then detail the maps, problem instances, and competing planners used to evaluate stages S2–S4, and define the performance metrics used to evaluate.
Training Procedure
We train a parametrized RL model $Q_{\theta}(s,a)$ using the DDQN algorithm(Van Hasselt, Guez, and Silver 2016), incorporating prioritized experience replay (PER) and $\varepsilon$‑greedy exploration, to enable an agent to navigate to a specified goal in the presence of both static and dynamic obstacles. Each dynamic obstacle possesses a hidden goal and follows precomputed trajectories, executing only valid moves; this allows for the simulation of online collision alerts during inference from the S3 stage. Training is performed for 30 00030\,000 episodes on an $11\times 11$ maze grid, with each episode capped at a maximum of $T_{\max}=50$ steps. The curriculum for environmental complexity is scheduled as follows: during the first 500 episodes, static‑obstacle density $\rho_{s}=0.10$ and dynamic‑obstacle count $n_{d}=0$; from episode 500 to 2999, $\rho_{s}=0.10$ and $n_{d}=1$; from episode 3000 to 5999, $\rho_{s}=0.20$ and $n_{d}=2$; and for episode 6000 onward, $\rho_{s}=0.30$ and $n_{d}=4$.
Input observations are normalized to zero mean and unit variance before being fed to the network. Transitions $(s_{t},a_{t},r_{t},s_{t+1})$ are stored in a prioritized experience replay buffer of size $10^{6}$. At each learning step, a mini‑batch of $128$ transitions is sampled. The agent selects actions using an $\varepsilon$‑greedy policy, where $\varepsilon$ decays from $\varepsilon_{0}=1.0$ to $0.01$ over the course of training according to $\varepsilon_{t+1}=\max(0.01,0.999\,\varepsilon_{t})$. Crucially, during both exploration and exploitation (e.g., when calculating the target Q-value), only actions $a_{t}$ deemed valid by an action mask $m_{t}\in\{0,1\}^{|\mathcal{A}|}$ are considered for efficient exploration (Damani et al. 2021b).
The Q-network parameters $\theta$ are updated via gradient descent using the Adam optimizer with a learning rate $\alpha=3\times 10^{-4}$. The discount factor is $\gamma=0.97$. A separate target network, with parameters $\theta^{-}$, provides stable targets and is updated by copying $\theta$ every 300 steps. Updates minimize the prioritized, importance‑weighted Bellman error. The target value $y_{t}$ and TD-error $\delta_{t}$ are computed as: $y_{t}=r_{t}+\gamma\,Q_{\theta^{-}}\bigl(s_{t+1},\arg\max_{a^{\prime}}Q_{\theta}(s_{t+1},a^{\prime})\text{ s.t. }m_{t}(a^{\prime})=1\bigr),$ $\delta_{t}=y_{t}-Q_{\theta}(s_{t},a_{t}),$ The loss function is then calculated as $L(\theta)\;=\;\mathbb{E}_{i\sim p(i)}\bigl[w_{i}\,\delta_{i}^{2}\bigr],$ where the sampling probability $p(i)\propto|\delta_{i}|^{\alpha}$ ($\alpha$ is a hyperparameter for PER) and $w_{i}$ are importance‑sampling weights that correct for the bias introduced by prioritized sampling.
In addition to the DDQN model, a Proximal Policy Optimization (PPO) based model was also trained; details regarding its architecture (including the one used for DDQN model) and training procedure, along with comparative performance, are provided in the Appendix Section A.3. The DDQN model was selected due to its superior performance over the PPO model - results are presented in Table 3.
Evaluation Scenarios and Baselines
We evaluate our framework’s performance and scalability across three distinct grid configurations, representing two primary map types: maze and warehouse. All evaluations use ten randomly generated problem instances per configuration. For the maze map type, we test on an $11\times 11$ grid with approximately 45% static obstacles (5 to 20 agents) and a larger $21\times 21$ grid with approximately 35% static obstacles (32, 64, and 96 agents). For the warehouse map type, we utilize a $25\times 25$ grid layout with approximately 24% static obstacles as shelves (32, 64, and 96 agents). Our proposed hybrid framework is evaluated through two variants, Alert-BFS and Alert-A*, which are named based on their RL policy rollout strategy (Breadth-First Search or weighted A-star search, respectively). The impact of our specific replanning strategies is further analyzed in an ablation study in Appendix Table 5. These are compared against search-based planners, Conflict-Based Search (CBS) (Sharon et al. 2015) and Improved CBS (ICBS) (Boyarski et al. 2015), and leading learning-based online planners PRIMAL (Damani et al. 2021a) and SCRIMP (Wang et al. 2023) with the MAPF formulation of disappear at goal. While our framework variants operate within our proposed four-stage pipeline (S1–S4), the baseline methods are executed as standalone algorithms. Per-instance time limits are 50 seconds for $11\times 11$ mazes, 30 minutes for $21\times 21$ mazes, and 1 hour for $25\times 25$ warehouse scenarios. For online planners, the step limits are 256 for the $11\times 11$ maze, and 512 for both the $21\times 21$ maze and the $25\times 25$ warehouse, as detailed in Table 1. Additional results on various other map configurations are provided in the Appendix Section A.4.
Performance Criteria
We evaluate our MAPF framework using standard metrics. Success Rate (SR) is the proportion of instances solved within defined time limits. For successful instances, Makespan (MS) measures the time until the last agent reaches its goal. The number of Collisions (CO) measures the total count of path disagreements detected and successfully resolved by the framework to produce the final collision-free solution. A non-zero CO in a successful run reflects the planner’s total conflict resolution workload. Finally, total Time (T) measures the computational cost until a solution is found or the time limit is exceeded. Learning-based methods are evaluated by the average number of steps taken in the environment.
5 Results and Discussion
Here we present an empirical study, evaluating our central hypothesis: a MAPF framework with strategically reduced information sharing can achieve robust performance. We address our RQs concerning minimal information needs and comparative performance against approaches with different information paradigms using the experimental results presented in Table 1.
(RQ1): Given the observation constraints of a decentralized setup, can an effective MAPF algorithm be created with one agent knowing nothing about other agents? If not, what information must it need at a minimum? Ans. No—if agents have zero knowledge of one another, MAPF fails, losing completeness guarantees; but solutions are possible with targeted, minimal information sharing. Table 5 in the appendix presents comparative results for various inter-agent information sharing settings.
Now, we turn our attention to the strategically limited minimal information sharing case. In our framework, strategically informed alerts convey either (S4.1) static constraints (the conflicting cell) or (S4.2) slightly richer, yet still localized, dynamic constraints (the sub-path of only the directly conflicting agent(s) for a short horizon). Table 1 shows performance comparison of considered methodologies across 11x11 maze (a), 21x21 maze (b), and $25\times 25$ (c) warehouse maps, respectively. As shown in Table 1, the performance of our framework, where agents initially plan with only local information and receive minimal, targeted alerts from a central coordinator upon conflict, achieves high Success Rates (SR). Specifically, on the $11\times 11$ and $21\times 21$ maze grids, SRs are high (90%-100%). On the more structured $25\times 25$ warehouse map (Table 1(c)), Alert-BFS and Alert-A* also achieve 100% SR for 32 and 64 agents, with Alert-A* maintaining a 60% SR in the dense agent scenario (96 agents), demonstrating effectiveness across different environment types with this minimal information sharing strategy. This success with minimal inter-agent information contrasts sharply with alternatives. Centralized planners (CBS, ICBS) demand complete global knowledge of all agents’ paths, a requirement that, while enabling optimality, leads to their low SR due to the combinatorial explosion in conflict resolution within practical time limits. In contrast, distributed learning frameworks (PRIMAL & SCRIMP) often operate under the assumption that each planning agent possesses full observation, of other agents and their goals, in a local field of view. While this localized observation can facilitate decentralized decision-making, it implies a continuous and potentially substantial information gathering (e.g., via onboard cameras or LiDAR sensors) and processing requirement for each agent.
The success of our method across maze and warehouse environments underscores that this on-demand information is sufficient for high SR, thereby avoiding the continuous, dense local information exchange. Our quantitative analysis, detailed in the Appendix A.1, shows that this on-demand alert mechanism reduces the total information load by $\sim$93% compared to a continuous-observation distributed paradigm in this scenario.
Table 1: Comparison of performance metrics for different methods on an $11\times 11$ grid (5–20 agents, $\sim$45% obstacles), a $21\times 21$ grid (32, 64, and 96 agents, $\sim$35% obstacles), and a $25\times 25$ warehouse grid (32, 64, 96 agents, $\sim$24% obstacles). Results are averaged over 10 problem instances per configuration. Key metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T), with Time averaged over all trials. For learning-based methods, the average steps taken while solving the problem instances are presented in brackets along with average wall-clock time taken.
(a) 11x11 maze with 50 sec. time limit per problem instance ($256$ steps for learning-based methods)
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (s) ($\downarrow$)
Methods 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20
Alert-BFS 100.00% 100.00% 98.00% 17.8 23.6 27.4 9 53 191 1.1 4.2 11.8
Alert-A* 100.00% 96.00% 82.00% 16.1 20.7 24.0 10 59 173 3.3 12.0 25.8
CBS 85.00% 34.00% 2.00% 13.8 14.3 17.0 559 8785 3895 8.3 30.7 35.3
ICBS 85.00% 26.00% 10.00% 14.0 14.8 12.8 885 3486 1466 5.9 25.1 28.6
PRIMAL 61.67% 48.00% 28.00% 71.7 89.9 104.8 - - - 1.4 (142) 3.4 (176) 6.1 (214)
SCRIMP 95.00% 70.00% 60.00% 19.3 28.6 38.8 1 2 8 0.4 (31) 1.6 (97) 3.9 (124)
(b) 21x21 maze with 30 min. time limit per problem instance ($512$ steps for learning-based methods)
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
Alert-BFS 100.00% 90.00% 0.00% 86.6 131.4 - 528 18909 - 2.7 23.9 30.0
Alert-A* 100.00% 80.00% 0.00% 41.4 72.0 - 457 11379 - 3.3 23.5 30.0
CBS 0.00% 0.00% 0.00% - - - - - - 30.0 30.0 30.0
ICBS 0.00% 0.00% 0.00% - - - - - - 30.0 30.0 30.0
PRIMAL 0.00% 0.00% 0.00% - - - - - - 0.8 (512) 2.1 (512) 1.8 (512)
SCRIMP 100.00% 20.00% 10.00% 65.9 135.0 94.0 15 220 357 0.1 (66) 1.0 (437) 2.4 (470)
(c) 25x25 warehouse with 1 hour time limit per problem instance ($512$ steps for learning-based methods)
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
Alert-BFS 100.00% 100.00% 10.00% 94.3 125.9 96.0 237 6379 26553 5.2 28.0 59.5
Alert-A* 100.00% 100.00% 60.00% 40.5 43.7 50.8 211 3129 16277 4.7 19.2 51.8
CBS 0.00% 0.00% 0.00% - - - - - - 60.0 60.0 60.0
ICBS 10.00% 0.00% 0.00% 38.0 - - 123107 - - 57.9 60.0 60.0
PRIMAL 0.00% 0.00% 0.00% - - - - - - 1.0 (512) 3.1 (512) 6.6 (512)
SCRIMP 100.00% 100.00% 90.00% 60.2 49.0 101.7 2 14 161 0.1 (60) 0.3 (49) 1.1 (143)
(RQ2): How does the proposed hybrid method compare to leading alternative search- and learning-based approaches in terms of performance, solution quality, and scalability? Ans. Our hybrid framework exhibits a compelling balance of performance and scalability with distinct trade-offs against established search-based and learning-based methods, as detailed in Table 1.
Our Python-based framework demonstrates higher SR and scalability compared to search-based optimal solvers. CBS/ICBS (C++ based) shows constantly lower SR with increasing environment complexity, failing to solve any instances on the $21\times 21$ maze (for 32+ agents) and the $25\times 25$ warehouse map. While CBS/ICBS achieves optimal MS when they solve (Table 1(a)), their reported Collision (CO) counts (Table 1(a)) for such cases are very high; our methods report significantly fewer collisions when they succeed on these smaller maps.
PRIMAL exhibits low SR across all tested configurations, often hitting step limits without success. SCRIMP is a strong learning-based performer with fast inference times and good MS. On the $11\times 11$ and $21\times 21$ mazes, its SR is comparable or better than our methods in some low-density cases, but can decline in denser scenarios ($21\times 21$, 64 agents: SCRIMP 20% SR vs. Alert-BFS 90% SR). On the $25\times 25$ warehouse map, SCRIMP maintains 100% SR for 32 and 64 agents, and achieves 90% SR for 96 agents, outperforming Alert-BFS (10% SR) and Alert-A* (60% SR) at this highest agent count in terms of success. However, SCRIMP also demonstrates significantly lower CO values than our methods across all successful warehouse instances attributing to its efficient multi-agent training. A critical factor for online planners is the offline training investment. For instance, PRIMAL’s extensive training on varied map sizes involved approximately 20 days on significant computational resources (Damani et al. 2021a). SCRIMP, with its sophisticated transformer-based architecture and PPO training, also implies a substantial training regimen across diverse configurations to achieve its reported performance (Wang et al. 2023). In contrast, our RL planning model, trained only on $11\times 11$ maze grids, was developed with a comparatively modest regimen: training was conducted on a Mac Mini equipped with an Apple M4 Pro chip (14-core CPU, 20-core GPU, 16-core Neural Engine) and 48GB unified memory, running for approximately 6 to 8 hours. Its subsequent effective performance when tested on different map types (map and warehouse) and larger grid sizes (up to $25\times 25$) indicates strong generalization from this simpler training setup.
6 Conclusions
Our hybrid framework combines decentralized RL-based planning with only minimal, targeted coordinated alerts–static conflict cell flags or brief conflict tracks–to maintain better SR across various challenging scenarios. While it incurs more collisions (CO) and sometimes longer makespan (MS) compared to highly specialized methods like SCRIMP, and its SR declines in extremely dense scenarios, it often delivers better overall feasibility and scalability, in scenarios with increasing agent counts, than purely search‑based methods and PRIMAL by strategically sharing minimum inter-agent information.
Effective multi-agent coordination with reduced information exchange, as explored in this research, will lead to privacy-aware automated systems, enabling broader use in complex privacy-preserving applications like autonomous driving. Despite these promising results, our work has limitations. The current framework, optimized for disappear at target setting, may face challenges in stay at target behavior. Future work will investigate the minimal information necessary to effectively manage these stay at target scenarios, aiming to extend the framework’s applicability without compromising its core principle of sparse information exchange. Additionally, an ablation study of richer AlertMask encodings, like dynamic occupancy maps, could improve replanning efficiency.
Appendix A Appendix
Appendix Contents
A.1. A Quantitative Model of Information Load ........................................................................................................................................................................A.1
A.2. Literature review ........................................................................................................................................................................A.2
    Literature Categorization Table 2 ........................................................................................................................................................................2
A.3. Training Methods ........................................................................................................................................................................A.3
  A.3. PPO Training Procedure ........................................................................................................................................................................A.3
  A.3. Neural Network Architecture ........................................................................................................................................................................A.3
  A.3. Model Comparison ........................................................................................................................................................................A.3
    DDQN training performance - Figure 2 ........................................................................................................................................................................
    PPO training performance - Figure 3 ........................................................................................................................................................................
  A.3. Model Comparison ........................................................................................................................................................................A.3
    Table 3 ........................................................................................................................................................................
A.4. Additional Experiments ........................................................................................................................................................................A.4
  A.4. Agent Selection Policy Comparison ........................................................................................................................................................................A.4
    Table 4 ........................................................................................................................................................................4
  A.4. Inter-agent Information Sharing Comparison ........................................................................................................................................................................A.4
    Table 5 ........................................................................................................................................................................5
  A.4. Benchmark Dataset Evaluation ........................................................................................................................................................................A.4
    Table 6 ........................................................................................................................................................................6
A.1 A Quantitative Model of Information Load
To quantitatively support our claim of reduced information sharing, we developed a simplified model to estimate the ”Total Information Load” of different coordination paradigms. This model is based on a simple Information Unit (IU), defined as the data required to represent one agent’s state (e.g., its coordinates) at a single timestep.
Model Assumptions
Our calculation is based on the 11x11 maze scenario with 20 agents, using metrics from the successful Alert-BFS runs in Table 1(a).
  1. 1.
    Scenario Data: We use a 20-agent instance (N=20) solved with a makespan (T) of 27 timesteps and requiring the resolution of 191 conflicts (C).
  2. 2.
    Average Path Length (L): We assume an average initial path length of 25 steps per agent.
  3. 3.
    Agent Density (D): For a distributed method like PRIMAL with a 10x10 FoV on an 11x11 grid, each agent has near-global vision. Therefore, each agent continuously observes all other N-1 agents. We set the average density (D) to 19 neighbors.
  4. 4.
    Alert Size ($I_{\text{alert}}$): We assume each conflict alert sent by our coordinator is a minimal, static alert costing 1 IU.
Comparative Calculation
Distributed Method (e.g., PRIMAL)
The information load is the continuous sensing and processing of all visible neighbors by every agent at every timestep. The formula is $\text{Info}_{\text{Distributed}}=N\times T\times D$.
$\text{Info}_{\text{Dist.}}$ $=20\times 27\times 19$
$=\textbf{10,260 IU}$
Our Hybrid Method (Alert-X)
The load is the one-time path sharing in addition to all subsequent targeted alerts. The formula is $\text{Info}_{\text{Alert-X}}=(N\times L)+(C\times I_{\text{alert}})$.
$\text{Info}_{\text{Alert-X}}$ $=(20\times 25)+(191\times 1)$
$=500+191$
$=\textbf{691 IU}$
Conclusion
This analysis highlights the difference in information architecture. The distributed method requires each of its 20 agents to continuously sense and process a heavy stream of local data, amounting to a total load of 10,260 IU. In contrast, our hybrid method offloads this burden to a central coordinator, resulting in a total information load of only 691 IU.
This represents a $\sim$93% reduction in the total information load, quantifying the efficiency of our on-demand alert system. While our method has a central coordinator, the burden on each individual agent is drastically lower, as it does not require constant, high-bandwidth sensing of its environment.
A.2 Literature review
Conflict-Based Search (CBS)
S1 – Agent Planning: Decentralized: Each agent independently computes its path from the start to the goal using a single-agent pathfinding algorithm (e.g., $A^{*}$). This decentralized planning allows for efficient initial path computation without considering other agents.
S2 – Collision Detection: Centralized: After individual paths are planned, CBS centrally examines these paths to detect conflicts, such as two agents occupying the same location at the same time (vertex conflicts) or swapping positions simultaneously (edge conflicts). This centralized detection ensures systematic identification of all potential conflicts.
S3 – Collision Avoidance Policy: Centralized: Upon detecting a conflict, CBS resolves it by adding constraints to the agents’ paths. Specifically, it creates two new branches in a constraint tree, each imposing a restriction on one of the conflicting agents to avoid the conflict. This centralized policy ensures optimal conflict resolution by exploring different constraint combinations (Sharon et al. 2015).
S4 – Agent Replanning: Decentralized: With new constraints in place, only the agents affected by these constraints replan their paths. Each affected agent independently computes a new path that adheres to the added constraints, maintaining the decentralized nature of the replanning process.
Large Neighborhood Search (LNS)
S1 – Agent Planning: Centralized: LNS-based methods, such as MAPF-LNS2, begin with a centralized planning phase where initial paths for all agents are computed using a fast, suboptimal solver like Prioritized Planning (PP). This initial solution may contain conflicts but serves as a starting point for further refinement (Li et al. 2022).
S2 – Collision Detection: Centralized: The system centrally identifies conflicts (e.g., vertex or edge collisions) in the initial set of paths. This global analysis ensures that all potential conflicts are detected and can be addressed in subsequent steps.
S3 – Collision Avoidance Policy: Centralized: LNS employs a centralized strategy to resolve conflicts by selecting subsets of agents (the ”neighborhood”) involved in collisions and replanning their paths. Techniques like Safe Interval Path Planning with Soft Constraints (SIPPS) are used to minimize the number of conflicts during this replanning phase.
S4 – Agent Replanning: Centralized: The selected subset of agents undergoes centralized replanning to resolve conflicts, while the paths of other agents remain unchanged. This process iterates, with different neighborhoods selected in each iteration, until a conflict-free set of paths is achieved or a predefined time limit is reached.
PRIMAL and PRIMAL2
S1 – Agent Planning: Distributed: In both PRIMAL and PRIMAL2, each agent independently plans its path using policies learned through a combination of RL and IL. These policies are trained to enable agents to navigate towards their goals based on local observations without centralized coordination (Sartoretti et al. 2019; Damani et al. 2021b).
S2 – Collision Detection: Distributed: Agents detect potential collisions based on their local observations of the environment. They do not rely on a centralized system to identify conflicts but instead use their learned policies to anticipate and respond to nearby agents.
S3 – Collision Avoidance Policy: Distributed: Collision avoidance is handled through the agents’ learned behaviors. In PRIMAL2, enhancements such as improved observation types have been introduced to facilitate better implicit coordination among agents, especially in dense and structured environments.
S4 – Agent Replanning: Distributed: Agents continuously replan their paths in response to changes in their local environment. This reactive planning allows them to adapt to dynamic scenarios, such as new goal assignments in lifelong MAPF settings.
SCRIMP
S1 – Agent Planning: Distributed: Each agent independently plans its path using a policy learned through a combination of RL and IL. Agents rely on a small local FOV (as small as 3×3) and a transformer-based communication mechanism to share information with nearby agents, enabling them to make informed decisions despite limited local observations (Wang et al. 2023).
S2 – Collision Detection: Distributed: Agents detect potential collisions based on their local observations and the messages received from neighboring agents through the communication mechanism. This decentralized approach allows agents to anticipate and respond to nearby agents without centralized coordination.
S3 – Collision Avoidance Policy: Distributed: Collision avoidance is handled through the agents’ learned policies, which incorporate a state-value-based tie-breaking strategy. This strategy enables agents to resolve conflicts in symmetric situations by assigning priorities based on predicted long-term collective benefits and distances to goals.
S4 – Agent Replanning: Distributed: Agents continuously replan their paths in response to changes in their local environment, leveraging intrinsic rewards to encourage exploration and mitigate the long-term credit assignment problem. This decentralized replanning allows agents to adapt to dynamic scenarios effectively.
Learn to Follow (FOLLOWER)
S1 – Agent Planning: Decentralized: Each agent independently plans its path to the assigned goal using a heuristic search algorithm (e.g., A*). To mitigate congestion, the planner incorporates a heatmap-based cost function that penalizes frequently occupied areas, encouraging agents to choose less crowded paths. A sub-goal (waypoint) is selected along the planned path to guide short-term movement (Skrynnik et al. 2024).
S2 – Collision Detection: Decentralized: Agents detect potential collisions based on their local observations. They do not rely on a centralized system to identify conflicts but instead use their learned policies to anticipate and respond to nearby agents.
S3 – Collision Avoidance Policy: Decentralized: Collision avoidance is handled through the agents’ learned behaviors. A neural network-based policy guides the agent toward its sub-goal while avoiding collisions. The policy is trained using reinforcement learning, leveraging local observations without requiring global state information or inter-agent communication.
S4 – Agent Replanning: Decentralized: Agents continuously replan their paths in response to changes in their local environment. This reactive planning allows them to adapt to dynamic scenarios, such as new goal assignments in lifelong MAPF settings.
LNS2+RL
S1 – Agent Planning: Centralized: LNS2+RL begins by centrally generating initial paths for all agents using a fast, suboptimal method like Prioritized Planning (PP). This initial solution may contain collisions but serves as a starting point for further refinement (Wang et al. 2025).
S2 – Collision Detection: Centralized: The system centrally identifies conflicts (e.g., vertex or edge collisions) in the initial set of paths. This global analysis ensures that all potential conflicts are detected and can be addressed in subsequent steps.
S3 – Collision Avoidance Policy: Hybrid: LNS2+RL employs a hybrid strategy for collision avoidance: Early Iterations: Utilizes a MARL policy to replan paths for subsets of agents involved in conflicts. This decentralized component allows agents to learn cooperative behaviors to avoid collisions. Later Iterations: Switches to a centralized PP algorithm for replanning, aiming to quickly resolve any remaining conflicts.
S4 – Agent Replanning: Hybrid: Replanning in LNS2+RL is conducted in a hybrid manner: Early Iterations: Selected subsets of agents undergo decentralized replanning using the MARL policy, promoting cooperative behavior. Later Iterations: Replanning shifts to a centralized approach using PP, focusing on efficiency and resolving any remaining conflicts.
Table 2: Related Literature Categorization into – S1: Agent Planning; S2: Collision Detection; S3: Collision avoidance policy; S4: Agent Replanning (SB: Search Based; LB: Learning Based)
Method S1 S2 S3 S4
Our* Decentralized Centralized Centralized Decentralized
(SB) CBS Decentralized Centralized Centralized Centralized
(SB) LNS Centralized Centralized Centralized Centralized
(LB) PRIMAL & PRIMAL-2 Distributed Distributed Distributed Distributed
(LB) SCRIMP Distributed Distributed Distributed Distributed
(LB) Learn to Follow Decentralized Decentralized Decentralized Decentralized
(LB) LNS2+RL Centralized Centralized Hybrid Hybrid
A.3 Training Methods
PPO Training Procedure
We train a single-agent navigation policy $\pi_{\theta}$ to reach a specified goal in the presence of both static and dynamic obstacles. Dynamic obstacles are each assigned a hidden goal and follow a precomputed trajectory, executing only valid moves; this setup allows us to generate online collision alerts during inference (cf. stage S3 of our fix–collisions algorithm).
Training is performed with Proximal Policy Optimization (PPO) (Schulman et al. 2017) on an $11\times 11$ maze grid for 30,00030{,}000 episodes. Throughout the first 15,00015{,}000 episodes, we linearly increase the static-obstacle density from $0\%$ to $30\%$ and the number of dynamic obstacles from 0 to 4.
At each time step $t$, we compute the discounted return
$R_{t}=\sum_{l=0}^{T-t}\gamma^{l}\,r_{t+l},$
and the advantage estimate using Generalized Advantage Estimation (GAE) (Schulman et al. 2015b):
$A_{t}=R_{t}-V_{\theta}(s_{t}),$
where $\gamma=0.95$ and $V_{\theta}(s_{t})$ is the value function.
The PPO surrogate objective is
$L_{\mathrm{PPO}}(\theta)=-\mathbb{E}_{t}\big[\min\big(r_{t}A_{t},\,\mathrm{clip}(r_{t},1-\varepsilon,1+\varepsilon)\,A_{t}\big)\big]$
$+c_{v}\,\mathbb{E}_{t}\big[\big(V_{\theta}(s_{t})-R_{t}\big)^{2}\big]-c_{e}\,\mathbb{E}_{t}\big[H\big(\pi_{\theta}(\cdot\mid s_{t})\big)\big],$
where $r_{t}=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\mathrm{old}}}(a_{t}\mid s_{t})}$, $\varepsilon=0.2$, $c_{v}=0.5$, and $c_{e}$ are linearly annealed from 0.05 down to 0.01 over the first 5,0005{,}000 episodes. PPO inherits many of the stability guarantees of trust-region methods (Schulman et al. 2015a).
To encourage the agent to distinguish valid from invalid moves, we add a binary cross-entropy loss
$L_{\mathrm{valid}}(\theta)=\mathbb{E}_{t}\big[\mathrm{BCE}\big(z_{t},m_{t}\big)\big],$
where $z_{t}\in\mathbb{R}^{|\mathcal{A}|}$ are the network logits, $m_{t}\in\{0,1\}^{|\mathcal{A}|}$ is the action-validity mask, and
$\mathrm{BCE}(z,m)=-\sum_{i=1}^{|\mathcal{A}|}\big[m_{i}\log\sigma(z_{i})+(1-m_{i})\log\big(1-\sigma(z_{i})\big)\big].$
The full objective is
$L(\theta)=L_{\mathrm{PPO}}(\theta)+\lambda_{\mathrm{valid}}\,L_{\mathrm{valid}}(\theta),$
where $\lambda_{\mathrm{valid}}$ is chosen empirically. During both training and inference, we sample only actions flagged as valid by $m_{t}$, which accelerates convergence and improves safety.
After training, the policy $\pi_{\theta}$ is used in stage S1 to generate initial trajectories and in stage S4 to replan whenever the centralized collision detector (stage S2) issues an alert.
Neural Network Architecture
ResNetDQN Network Architecture for DDQN Training
ResNetDQN is a residual‐network architecture for approximating the action‐value function $Q^{\pi}(s,a)$ in grid‐based environments trained using the DDQN algorithm (Van Hasselt, Guez, and Silver 2016). It combines a deep convolutional stem with residual blocks (He et al. 2016), early fusion of low‐dimensional features, hierarchical downsampling, and a late‐fusion MLP to output one Q‐value per action.
Network Overview
  1. 1.
    Input: $(B,6,H,W)$ tensor comprising four binary masks (obstacle, agent, goal, dynamic) plus normalized $x$- and $y$-coordinate channels, and a $(B,3)$ low‐dim vector of {direction, distance}.
  2. 2.
    Conv Stem & Residual Blocks: $3\times 3$ conv (6→32 channels, stride=1, pad=1) + BatchNorm + ReLU, then two ResidualBlock(32→32) with dilation rates 1 and 2.
  3. 3.
    Early Fusion: Project low‐dim (3→16), tile to $(B,16,H,W)$, concat with conv features → 48 channels, then $3\times 3$ conv (48→32) + BatchNorm + ReLU.
  4. 4.
    Downsampling Stage 1: $3\times 3$ conv (32→64, stride=2, pad=1) + BatchNorm + ReLU, followed by two ResidualBlock(64→64) with dilation rates 2 and 4.
  5. 5.
    Downsampling Stage 2: $3\times 3$ conv (64→128, stride=2, pad=1) + BatchNorm + ReLU, followed by two ResidualBlock(128→128) with dilation rates 4 and 1.
  6. 6.
    Global Pooling & Late Fusion: AdaptiveAvgPool2d → 128‐dim vector $\mathbf{u}$. In parallel, map low‐dim→256, map $\mathbf{u}\to 256$, concat →512 →256 via FC + ReLU.
  7. 7.
[…]
  8. 8.
    Initialization: All conv and linear weights: Xavier‐uniform; all biases: zero.
Implementation and Packages
  • torch, torch.nn, torch.nn.functional: define modules, layers, activations.
  • AdaptiveAvgPool2d, BatchNorm2d, Conv2d, Linear: core building blocks.
ResNet-based Actor–Critic Architecture for PPO Training
The PPOActorCritic model shares the same ResNet encoder as ResNetDQN (conv stem, residual blocks, fusion, downsampling, pooling) (He et al. 2016), producing a 256‐dim embedding. It splits into two GRU‐based heads for policy (actor) and value (critic), trained via PPO (Schulman et al. 2017).
Network Overview
  1. 1.
    Shared Encoder: Follows Steps 1–4 from ResNetDQN, yielding a 256‐dim hidden $\mathbf{h}_{\mathrm{shared}}$.
  2. 2.
[…]
  3. 3.
    Critic Head: GRUCell(256→256) updates hidden state $\mathbf{h}^{V}_{t}$; Linear(256→1) produces scalar state‐value estimate.
  4. 4.
    Initialization: Same Xavier‐uniform for all weights, zero biases.
Implementation and Packages
  • torch, torch.nn, torch.nn.functional: define encoder, GRUs, heads.
  • GRUCell, Linear: recurrent and output modules.
Training Results
Figure 2: The plot illustrates the training performance of the Double Deep Q-Network (DDQN) algorithm. Episode rewards (red), sample efficiency measured by rewards per total frames (purple), training loss (blue), and episode length (orange) are presented across episodes. Smoothed curves represent moving averages, enhancing the visibility of underlying performance trends.
Figure 3: The plot illustrates the training performance of the Proximal Policy Optimization (PPO) algorithm, capturing episode rewards (red). Episode rewards (red), sample efficiency measured by rewards per total frames (purple), training loss (blue), and episode length (orange) are presented across episodes. Smoothed curves represent moving averages, enhancing the visibility of underlying performance trends.
Model Comparison
Table 3: Performance comparison of reinforcement learning methods (BFS-DQN, BFS-PPO, A*-DQN, A*-PPO) across different grid environments and agent counts. Results are averaged over 10 problem instances per configuration. Evaluated metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T), with the Time metric averaged over all trials. Shows that the model trained using DDQN, selected for our main paper, yields better results over PPO model.
(a) 11x11 maze grid, $\sim$45% static obstacles (5–20 agents). Time in seconds.
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (s) ($\downarrow$)
Methods 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20
BFS-DDQN 100.00% 100.00% 98.00% 17.8 23.6 27.4 9 53 191 1.1 4.2 11.8
BFS-PPO 100.00% 94.00% 74.00% 22.1 28.6 28.6 9 48 125 4.9 16.5 31.8
A*-DDQN 100.00% 96.00% 82.00% 16.1 20.7 24.0 10 59 173 3.3 12.0 25.8
A*-PPO 100.00% 96.00% 74.00% 16.6 21.4 25.2 10 60 171 3.9 14.4 30.4
(b) 21x21 maze grid, $\sim$35% static obstacles (32, 64, and 96 agents). Time in minutes.
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
BFS-DDQN 100.00% 90.00% 0.00% 86.6 131.4 - 528 18909 - 2.7 23.9 33.5
BFS-PPO 100.00% 0.00% 0.00% 87.6 - - 463 - - 11.2 30.2 31.1
A*-DDQN 100.00% 80.00% 0.00% 41.4 72.0 - 457 11379 - 3.3 23.5 33.4
A*-PPO 100.00% 0.00% 0.00% 39.6 78.3 - 407 10761 - 6.2 27.0 30.2
(c) 25x25 warehouse grid, $\sim$24% static obstacles (32, 64, and 96 agents). Time in minutes.
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
BFS-DDQN 100.00% 100.00% 10.00% 94.3 125.9 96.0 237 6379 26553 5.2 28.0 59.5
BFS-PPO 100.00% 40.00% 0.00% 93.6 130.0 - 231 3740 - 15.2 57.3 60.2
A*-DDQN 100.00% 100.00% 60.00% 40.5 43.7 50.8 211 3129 16277 4.7 19.2 51.8
A*-PPO 100.00% 100.00% 40.00% 41.6 42.4 51.5 179 3117 17382 5.7 30.0 55.4
A.4 Additional Experiments
We present a model comparison trained using DDQN and PPO reinforcement learning algorithms, Table 3, in Section A.3. This section details further empirical evaluations and an ablation study of our framework.
An ablation study of different agent selection policies is detailed in Table 4. We have considered the information sharing setting/replanning strategy as Static & Dynamic Obstacle Replanning.
Furthermore, a comparison of replanning settings for static and dynamic obstacles is provided in Table 5, which reflects different information sharing settings as detailed in our main paper. Here, we considered the agent selection policy as Fewest Future Collisions (FFC).
We also present extended analyses on MovingAI lab’s MAPF benchmark dataset (Stern et al. 2019b) – random map $32\times 32-10$ environment with various agent counts [8, 16, 32, 64, 96], results presented in Section A.4.
Agent Selection Policy Comparison
Table 4: Performance comparison of Agent Selection Policies (Random, Farthest, Fewest Future Collisions - FFC) for Alert-BFS and Alert-A* methods across different grid environments and agent counts. Metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T). Results are averaged over 10 problem instances per configuration. Best performance across policies for each method and agent group is bolded. Using Static & Dynamic obstacle replanning. Shows that, among the choices [random, farthest & fewest future collisions (FFC)], the FFC policy, selected for our main paper, yields best results across the different map configurations.
(a) 11x11 maze, $\sim$45% static obstacles (5–20 agents). Time in seconds (s).
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (s) ($\downarrow$)
Methods 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20
Policy: Random Agent Selection
Alert-BFS 98.33% 98.00% 82.00% 20.5 24.6 32.7 20 108 517 2.0 5.8 22.6
Alert-A* 98.33% 96.00% 50.00% 17.8 23.6 24.3 16 122 259 4.3 15.7 35.6
Policy: Farthest Agent Selection
Alert-BFS 100.00% 94.00% 90.00% 19.8 24.6 30.3 14 91 387 1.2 6.9 17.8
Alert-A* 98.33% 84.00% 62.00% 18.9 23.3 26.5 14 96 224 4.6 18.3 35.1
Policy: Fewest Future Collisions (FFC)
Alert-BFS 100.00% 100.00% 98.00% 17.8 23.6 27.4 9 53 191 1.1 4.2 11.8
Alert-A* 100.00% 96.00% 82.00% 16.1 20.7 24.0 10 59 173 3.3 12.0 25.8
(b) 21x21 maze grid, $\sim$35% static obstacles (32, 64, and 96 agents). Time in minutes (min).
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
Policy: Random Agent Selection
Alert-BFS 100.00% 0.00% 0.00% 99.2 - - 1670 - - 4.1 30.1 30.4
Alert-A* 100.00% 0.00% 0.00% 48.8 - - 931 - - 6.1 30.2 30.2
Policy: Farthest Agent Selection
Alert-BFS 100.00% 0.00% 0.00% 93.5 - - 1216 - - 3.1 30.2 30.2
Alert-A* 100.00% 10.00% 0.00% 47.5 64.0 - 697 14654 - 5.3 29.8 30.3
Policy: Fewest Future Collisions (FFC)
Alert-BFS 100.00% 90.00% 0.00% 86.6 131.4 - 528 18909 - 2.7 23.9 30.5
Alert-A* 100.00% 80.00% 0.00% 41.4 72.0 - 457 11379 - 3.3 23.5 30.4
(c) 25x25 warehouse grid, $\sim$24% static obstacles (32, 64, and 96 agents). Time in minutes (min).
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
Policy: Random Agent Selection
Alert-BFS 100.00% 20.00% 0.00% 119.0 139.0 - 722 23853 - 12.4 57.9 60.1
Alert-A* 100.00% 100.00% 0.00% 42.4 54.6 - 302 5788 - 6.0 37.9 60.1
Policy: Farthest Agent Selection
Alert-BFS 100.00% 60.00% 0.00% 83.4 126.0 - 455 9836 - 8.5 47.6 60.8
Alert-A* 100.00% 100.00% 0.00% 45.0 59.4 - 268 4644 - 6.4 34.9 60.1
Policy: Fewest Future Collisions (FFC)
Alert-BFS 100.00% 100.00% 10.00% 94.3 125.9 96.0 237 6379 26553 5.2 28.0 59.5
Alert-A* 100.00% 100.00% 60.00% 40.5 43.7 50.8 211 3129 16277 4.7 19.2 51.8
Inter-agent Information Sharing Comparison
Table 5: Comparison of replanning settings (Only Static, Only Dynamic, Static & Dynamic Obstacles) for Alert-BFS and Alert-A* methods across different grid environments and agent counts. Metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T). Results are averaged over 10 problem instances per configuration. Best performance across settings for each method and agent group is bolded. Using FFC agent selection policy. Shows that, among the replanning strategies, Static + Dynamic obstacle replanning strategy (selected for our main paper), yields best results across the different map configurations.
(a) 11x11 maze grid, $\sim$45% static obstacles (5–20 agents). Time in seconds (s).
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (s) ($\downarrow$)
Methods 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20
Setting 1: Only Static Obstacle Replanning (No inter-agent information)
Alert-BFS 95.00% 92.00% 84.00% 17.5 23.0 25.4 8 51 205 1.4 4.6 14.2
Alert-A* 95.00% 90.00% 78.00% 15.6 20.3 23.7 8 55 177 3.2 11.6 26.0
Setting 2: Only Dynamic Obstacle Replanning
Alert-BFS 63.33% 24.00% 8.00% 19.3 21.0 26.8 16 89 171 4.9 17.9 34.7
Alert-A* 50.00% 12.00% 2.00% 14.0 13.8 15.0 6 25 49 22.9 41.4 52.1
Setting 3: Static & Dynamic Obstacle Replanning
Alert-BFS 100.00% 100.00% 98.00% 17.8 23.6 27.4 9 53 191 1.1 4.2 11.8
Alert-A* 100.00% 96.00% 82.00% 16.1 20.7 24.0 10 59 173 3.3 12.0 25.8
(b) 21x21 maze grid, $\sim$35% static obstacles (32, 64, and 96 agents). Time in minutes (min).
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
Setting 1: Only Static Obstacle Replanning (No inter-agent information)
Alert-BFS 100.00% 10.00% 0.00% 86.6 140.0 - 528 12476 - 5.8 30.0 30.0
Alert-A* 100.00% 20.00% 0.00% 41.4 59.0 - 457 9310 - 6.7 30.0 30.0
Setting 2: Only Dynamic Obstacle Replanning
Alert-BFS 0.00% 0.00% 0.00% - - - - - - 30.0 30.0 30.0
Alert-A* 0.00% 0.00% 0.00% - - - - - - 30.0 30.0 30.0
Setting 3: Static & Dynamic Obstacle Replanning
Alert-BFS 100.00% 90.00% 0.00% 86.6 131.4 - 528 18909 - 2.7 23.9 30.0
Alert-A* 100.00% 80.00% 0.00% 41.4 72.0 - 457 11379 - 3.3 23.5 30.0
(c) 25x25 warehouse grid, $\sim$24% static obstacles (32, 64, and 96 agents). Time in minutes (min).
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
Setting 1: Only Static Obstacle Replanning (No inter-agent information)
Alert-BFS 100.00% 100.00% 0.00% 95.8 117.4 - 230 6097 - 10.0 40.0 60.2
Alert-A* 100.00% 100.00% 60.00% 42.2 41.6 56.0 204 3258 16694 8.9 32.3 53.3
Setting 2: Only Dynamic Obstacle Replanning
Alert-BFS 0.00% 0.00% 0.00% - - - - - - 61.7 62.4 60.6
Alert-A* 0.00% 0.00% 0.00% - - - - - - 58.6 61.6 61.8
Setting 3: Static & Dynamic Obstacle Replanning
Alert-BFS 100.00% 100.00% 10.00% 94.3 125.9 96.0 237 6379 26553 5.2 28.0 59.5
Alert-A* 100.00% 100.00% 60.00% 40.5 43.7 50.8 211 3129 16277 4.7 19.2 51.8
Benchmark Dataset Evaluation
Map Configuration: $32\times 32-10$ Random Map
The random map $32\times 32-10$ is obtained from the Moving AI Lab’s MAPF benchmark dataset (Stern et al. 2019b). Agent configurations tested include scenarios with 8, 16, 32, 64, and 96 agents. Each configuration was tested across 10 different problem instances (unique start and goal locations for all agents). The detailed performance metrics, including success rates, makespan, collisions, and computation time, are presented in Table 6.
Table 6: Comparison of performance metrics for different methods on a $32\times 32$ random map. Results are averaged over 10 problem instances per configuration and presented in two subtables based on agent counts. Key metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T(min)), with Time averaged over all trials. For learning-based methods like PRIMAL and SCRIMP, the average steps taken (episode length) are presented in brackets along with average wall-clock time taken. For learning based methods, the environment step limit is set to 512. Best performing values for each metric per agent count are highlighted in bold.
(a) 32x32 random map (8, 16, and 32 agents). The time limit per problem instance is 1 hour. For learning based methods, the environment step limit is set to 512.
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 8 16 32 8 16 32 8 16 32 8 16 32
Alert-BFS 100.00% 100.00% 100.00% 62.5 76.2 104.9 2 9 64 0.2 0.6 2.4
Alert-A* 100.00% 100.00% 100.00% 42.7 45.5 48.1 2 8 68 0.2 0.9 2.5
CBS 100.00% 100.00% 90.00% 40.2 43.6 47.0 1 6 35079 0.01 0.01 13.0
ICBS 100.00% 100.00% 90.00% 40.2 43.6 47.0 1 6 15438 0.01 0.01 10.7
PRIMAL 0.00% 0.00% 0.00% - - - - - - 0.1 (512) 0.3 (512) 1.0 (512)
SCRIMP 100.00% 100.00% 100.00% 39.4 43.3 47.7 0 0 0 0.01 (39) 0.01 (43) 0.03 (48)
(b) 32x32 random map (64 and 96 agents). The time limit per problem instance is 1 hour. For learning based methods, the environment step limit is set to 512.
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 64 96 64 96 64 96 64 96
Alert-BFS 100.00% 100.00% 136.5 154.4 1194 7079 12.9 33.8
Alert-A* 100.00% 100.00% 53.2 52.0 1158 5912 13.2 30.7
CBS 0.00% 0.00% - - - - 60.1 60.1
ICBS 0.00% 0.00% - - - - 60.3 60.2
PRIMAL 0.00% 0.00% - - - - 3.6 (512) 8.4 (512)
SCRIMP 100.00% 90.00% 52.1 55.8 2 3 0.2 (52) 0.5 (101)

### v2

@@
-Scalable Multi-Agent Path Finding using Collision-Aware Dynamic Alert Mask and a Hybrid Execution Strategy
+Towards Information-Optimized Multi-Agent Path Finding: A Hybrid Framework with Reduced Inter-Agent Information Sharing
 Abstract
 Multi-agent pathfinding (MAPF) remains a critical problem in robotics and autonomous systems, where agents must navigate shared spaces efficiently while avoiding conflicts.
-Traditional centralized algorithms that have global information, such as Conflict-Based Search (CBS), provide high-quality solutions but become computationally expensive in large-scale scenarios due to the combinatorial explosion of conflicts that need resolution.
+Traditional centralized algorithms with global information provide high-quality solutions but scale poorly in large-scale scenarios due to the combinatorial explosion of conflicts.
 Conversely, distributed approaches that have local information, particularly learning-based methods, offer better scalability by operating with relaxed information availability, yet often at the cost of solution quality.
-To address these limitations, we propose a hybrid framework that combines decentralized path planning with a lightweight centralized coordinator.
-Our framework leverages reinforcement learning (RL) for decentralized planning, enabling agents to adapt their planning based on minimal, targeted alerts–such as static conflict-cell flags or brief conflict tracks–that are dynamically shared information from the central coordinator for effective conflict resolution.
-We empirically study the effect of the information available to an agent on its planning performance.
-Our approach reduces the inter-agent information sharing compared to fully centralized and distributed methods, while still consistently finding feasible, collision-free solutions–even in large-scale scenarios having higher agent counts.
+In realistic deployments, information is a constrained resource: broadcasting full agent states and goals can raise privacy concerns, strain limited bandwidth, and require extra sensing and communication hardware, increasing cost and energy use.
+We focus on the core question of how MAPF can be solved with minimal inter-agent information sharing while preserving solution feasibility.
+To this end, we present an information-centric formulation of the MAPF problem and introduce a hybrid framework, IO-MAPF, that integrates decentralized path planning with a lightweight centralized coordinator.
+In this framework, agents use reinforcement learning (RL) to plan independently, while the central coordinator provides minimal, targeted signals, such as static conflict-cell indicators or short conflict trajectories, that are dynamically shared to support efficient conflict resolution.
+We introduce an Information Units (IU) metric to quantify information use and show that our alert-driven design achieves $2\times$ to $23\times$ reduction in information sharing, compared to the state-of-the-art algorithms, while maintaining high success rates, demonstrating that reliable MAPF is achievable under strongly information-restricted, privacy-preserving conditions.
+We demonstrate the effectiveness of our algorithm using both simulation and hardware experiments.
 1 Introduction
-Multi-Agent Path Finding (MAPF) addresses the fundamental challenge of computing collision-free trajectories for multiple agents navigating a shared environment.
-Its effective solving is crucial for deploying a wide array of multi-agent systems, from automated warehouses and robotic swarms to autonomous vehicle coordination (Sharon et al.
-2015).
-Despite its practical significance, MAPF is computationally demanding, classified as NP-hard in its general form, which can render traditional centralized search methods intractable as the number of agents or the complexity of the environment increases (Ren et al.
-2025; Sartoretti et al.
-2019).
+Multi-Agent Path Finding (MAPF) addresses the fundamental problem of computing collision-free trajectories for multiple agents navigating a shared environment.
+Solving it effectively is crucial for a wide range of real-world applications, ranging from the deployment of automated robotic swarms in warehouses to performing autonomous vehicle coordination.
+Despite its practical significance, solving MAPF is computationally demanding, classified as NP-hard in its general form, which can render traditional centralized search methods intractable as the number of agents or the complexity of the environment increases (Ren et al.
+2025; Sartoretti and others 2019).
 Figure 1: An example MAPF problem and our four-stage planning pipeline.
@@
 Bottom Right: The four stages of our framework, from initial Path Planning (S1) to Collision Detection (S2), Resolution (S3), and Replanning (S4).
-Extensive research in MAPF has explored various coordination paradigms, each with different implications for information availability and system performance.
+The MAPF literature has explored various coordination paradigms, each with different implications for information availability and system performance.
 Centralized approaches, such as Conflict-Based Search (CBS) (Sharon et al.
 2015) and its efficiency-focused variant ICBS (Boyarski et al.
-2015), typically assume full observability of the environment and agent states, helping them generate optimal plans.
+2015), typically assume full knowledge of the environment and agent states and compute collision-free paths based on this global information available.
 However, as the number of agents or the size of the environment increases, joint planning and conflict resolution become computationally expensive.
-Additionally, requiring agents to share full information raises privacy concerns, limiting applicability in confidential settings motivating strategies that reduce reliance on global information.
+Additionally, requiring agents to share full information raises privacy concerns, limiting applicability in confidential settings and imposing additional compute requirements in real-world deployments, motivating strategies that reduce reliance on global information.
 Early decoupled methods like $M^{*}$ (Wagner and Choset 2011) begin with individual plans and only coordinate agents in conflict, offering scalability but often sacrificing optimality or completeness.
 More recently, Multi-Agent Reinforcement Learning (MARL) has emerged to address coordination and partial observability in dynamic settings.
-For example, PRIMAL (Sartoretti et al.
-2019) trains agents to plan using partial views, learning implicit coordination, while methods like FOLLOWER (Skrynnik et al.
+For example, PRIMAL (Sartoretti and others 2019) trains agents to plan using partial views, learning implicit coordination, while methods like FOLLOWER (Skrynnik et al.
 2024) reduce explicit communication by relying on global heuristic maps.
 These decentralized methods improve scalability but may reduce solution quality and leave conflicts unresolved due to limited information.
-To address these limitations, we propose a novel hybrid MAPF framework that combines decentralized planning with a lightweight centralized coordinator, focusing on minimizing inter-agent information sharing while maintaining solution feasibility.
-Our findings show that minimal, targeted alerts are sufficient, reducing the total information load by an estimated $\sim$93% compared to a continuous-observation distributed paradigm (see Appendix A.1).
+To address these limitations, we first formulate the MAPF problem from an information-centric perspective and with the aim of reducing information exchange under this setting, we propose Information-Centric MAPF (IC–MAPF), a novel hybrid MAPF framework that combines decentralized planning with a lightweight centralized coordinator, that minimizes inter-agent information sharing while maintaining solution feasibility.
+Our findings show that minimal, targeted alerts are sufficient, reducing the total information load by an estimated $2\times$ to $23\times$ compared to the state-of-the-art communication efficient algorithms (see Table 3).
 In our approach, agents primarily rely on decentralized, reinforcement learning-based neural network planners that operate on local observations (e.g., position coordinates), eliminating the need for global information as well as local egocentric maps.
@@
 This work’s contributions are centered on this selective coordination strategy.
-We introduce (1) a novel hybrid framework with this on-demand alert mechanism and provide (2) a comprehensive empirical evaluation demonstrating that our planning policy generalizes robustly from a simple training regimen to larger, more complex environments.
+We introduce (1) a information-centric MAPF formulation and identify the information utilized by various MAPF strategies, (2) IC–MAPF - a novel hybrid framework with an on-demand alert mechanism that substantially reduces the amount of global information an agent must access for successful planning, and (3) a new metric, Information Units (IU), to quantify the information available to each agent during planning.
 Our evaluation is guided by two key research questions: (RQ1) Given the observation constraints of a decentralized setup, can an effective MAPF algorithm be created with one agent knowing nothing about other agents?
@@
 2015).
-Common efficiency objectives include minimizing the makespan, $\max_{i}T_{i}$, minimizing the sum of individual completion times (sum-of-costs), $\sum_{i}T_{i}$, or minimizing the number of collisions.
-An important consideration in defining a MAPF problem instance arises from agents potentially reaching their targets at different time steps, thus requiring a clear definition of how an agent behaves after it has reached its target $g_{i}$ at time $T_{i}$, but while other agents may still be en route.
-Two common settings are used to address this scenario (Stern et al.
+Common efficiency objectives include minimizing the makespan, $\max_{i}T_{i}$, minimizing the sum of individual completion times (sum-of-costs), $\sum_{i}T_{i}$.
+An important modeling choice in MAPF concerns how agents are treated after reaching their goals at potentially different times.
+Two common settings are used (Stern et al.
 2019a).
-The first setting, typically referred to as stay at target, considers that an agent, upon reaching its goal, remains at that vertex until all other agents in the team have also reached their respective targets.
-This waiting agent continues to occupy its target vertex, potentially causing conflicts if another agent’s plan requires traversing or occupying that vertex at any time $t\geq T_{i}$ up to the completion time of the entire solution.
-The second setting, disappear at target, assumes an agent is effectively removed from the environment immediately upon reaching its target.
-Consequently, its path concludes at time $T_{i}$, and it poses no further collision risk nor occupies any vertex for $t>T_{i}$.
-In our work, we adopt the second setting, disappear at target.
-This setting is particularly relevant for applications like drone delivery, where an agent’s task is considered complete upon arrival at a target location.
+In the stay-at-target setting, an agent remains at its goal vertex after arrival and continues to occupy that location until all agents have finished, potentially creating conflicts for others.
+In contrast, the disappear-at-target setting removes an agent from the environment immediately upon goal arrival, eliminating any further interactions.
+In this work, we adopt the stay-at-target setting, which is appropriate for persistent agents operating in shared environments.
 2.2 Coordination Paradigms
@@
 In decentralized MAPF, as defined in this work, inter-agent communication during execution is eliminated.
-Agents plan and act entirely independently, relying only on their own local information.
+Agents plan and act entirely independently, relying only on their own path information.
 Finally, hybrid frameworks, like the one we propose, combine these elements.
@@
 2.3 Literature Review
-Search-based Methods: Research on MAPF has evolved from centralized to more scalable methods.
-Early advancements include $M^{*}$, which dynamically couples agent searches upon conflict for optimality and completeness (Wagner and Choset 2011), and Increasing Cost Tree Search (ICTS), which employs cost allocation and then checks for conflict-free solutions from sets of single-agent paths, yielding significant speedups over joint-space search (Sharon et al.
+Search-based Methods: Classical MAPF research progressed from centralized optimal solvers to more scalable search frameworks.
+Early methods such as $M^{*}$ dynamically couple agents only when conflicts arise (Wagner and Choset 2011), while ICTS allocates costs to agents and searches for conflict-free combinations of single-agent paths, reducing joint-space complexity (Sharon et al.
 2013).
-A pivotal development was CBS, where agents independently plan paths using $A^{*}$, and a centralized detector resolves conflicts by branching with new constraints for involved agents, ensuring optimality (Sharon et al.
+A major milestone was CBS, which plans independently for each agent using $A^{*}$ and resolves conflicts by branching on constraints, guaranteeing optimality (Sharon et al.
 2015).
-To enhance scalability while preserving optimality, variants such as ICBS integrate cardinality heuristics and meta-agent merging to accelerate convergence (Boyarski et al.
+Its variants, including ICBS, further accelerate search through cardinality heuristics and meta-agent merging (Boyarski et al.
 2015).
-For very large instances, suboptimal repair strategies like Large Neighborhood Search (LNS) generate initial solutions (often via prioritized planning) and iteratively repair colliding agent subsets under a centralized control loop (Li et al.
+For large-scale problems, suboptimal approaches like LNS generate an initial solution (often via prioritized planning) and iteratively repair conflicting subsets (Li et al.
 2022).
-However, these methods face scalability challenges as the number of agents increases and often depend on extensive global information for coordination.
-Learning-based Methods: In parallel, learning‐based approaches address partial observability and communication constraints.
-PRIMAL and PRIMAL2 train decentralized policies via a combination of imitation learning from an expert centralized planner and reinforcement learning (PRIMAL2 enhancing local observations for better coordination), effectively implementing a centralized training and decentralized execution (CTDE) paradigm (Sartoretti et al.
-2019; Damani et al.
-2021b).
-SCRIMP introduces a transformer‐based communication module enabling agents with limited fields of view to coordinate effectively, achieving decentralized coordination without a central execution-time coordinator (Wang et al.
+PIBT (Okumura et al.
+2022) introduces adaptive prioritization for local agent movements, and LaCAM (Okumura 2023) combines low-level constraint extraction with high-level location-sequence search.
+Despite their efficiency gains, these methods still face scalability limits with growing agent counts and rely heavily on global information.
+Learning-based Methods: Learning-based MAPF methods focus on handling partial observability and limited communication through decentralized policies trained under the centralized training and decentralized execution (CTDE) paradigm.
+PRIMAL and PRIMAL2 combine imitation learning from an expert centralized solver with reinforcement learning (PRIMAL2 further improving local observations), enabling scalable decentralized execution (Sartoretti and others 2019; Damani et al.
+2021).
+SCRIMP employs a transformer-based communication module that supports coordination among agents with restricted field-of-view (FOV), allowing decentralized conflict avoidance without a central coordinator (Wang et al.
 2023).
-To function in real‐world settings, each agent must carry onboard sensing, e.g., RGB‐D cameras or LiDAR scanners, and implement an inter-agent communication protocol for goal disclosure, which introduces additional computational overhead and raises privacy concerns.
-While offering adaptability and decentralized execution, learning-based methods can require significant training data, may struggle with generalization, and often provide weaker guarantees on solution quality or completeness when relying purely on local information.
-Hybrid Methods: Hybrid planning‐learning paradigms seek the strengths of both worlds.
-Authors in (Skrynnik et al.
-2024) introduce FOLLOWER, which involves each agent using a congestion‐aware $A^{*}$ planner to determine its subgoals, then executing a fully decentralized RL policy for lifelong replanning.
-More recently, hybrid frameworks like LNS2+RL, introduced by Li et al., have emerged, augmenting LNS by employing MARL for early, localized conflict resolution, then relying on efficient search-based planning for final refinement (Wang et al.
-2025).
-These existing hybrid approaches, while innovative, can involve complex integrations and may still face trade-offs between optimality, scalability, or specific information dependencies (like global congestion maps (Skrynnik et al.
-2024)).
-To enable a structured comparison of diverse MAPF techniques (search-based, learning-based, and hybrid), we adopt the four‑stage MAPF pipeline: S1 (Agent Planning), S2 (Collision Detection), S3 (Collision Avoidance Policy), and S4 (Agent Replanning).
-Within this structure, CBS uses decentralized A* search for S1/S4 and centralized high level solver for collision resolution in S2/S3 for optimality (Sharon et al.
-2015).
-LNS employs a centralized loop for all four stages, managing path generation, detection, repair, and subset replanning, thereby trading optimality for scalability (Li et al.
-2022).
-Learning-based planners like PRIMAL and SCRIMP typically use local observations for decentralized S1-S4 execution, coordinating without a central controller (Sartoretti et al.
-2019; Wang et al.
-2023); however, practical deployments can require sensors (eg., camera, LiDAR etc.) and communication, adding computational overhead and increase sensing cost.
-Hybrid methods such as FOLLOWER use global congestion maps with $A^{*}$ search for S1, then decentralized RL policy for S2-S4, often without further communication (Skrynnik et al.
-2024).
-LNS2+RL applies centralized LNS for S1-S3, while S4 uses a hybrid policy (MARL or prioritized planning) for subsets, balancing key trade-offs (Wang et al.
-2025).
-Refer to Table 2 in the Appendix Section A.2 for a detailed categorization of these methodologies.
-Together, these works highlight key limitations of existing MAPF paradigms: centralized methods face scalability and privacy concerns; purely learning-based approaches compromise solution quality; and current hybrid solutions often still depend on significant centralized communication or coordination overhead with higher sensing cost.
-The hybrid framework introduced in Section 3, which is the focus of our work, is specifically designed to mitigate these limitations.
-It strategically reduces inter-agent information sharing to maintain solution feasibility, thereby addressing the aforementioned information management challenges while also aiming to preclude the additional sensing costs and computational overhead common in learning-based methods.
+SYLPH (He et al.
+2025) enables agents to infer their Social Value Orientation (SVO), a measure of selfishness or altruism, based on situational context, and adapt their behavior toward the most influential agent in the system.
+ALPHA (He et al.
+2024) blends accurate proximal sensing with fuzzy global cues to guide agents beyond purely local reasoning, mitigating local myopia.
+SIGMA (Liao et al.
+2025) uses sheaf-theoretic local consensus to capture geometric cross-dependencies among agents.
+In real deployments, each agent must carry onboard sensors (e.g., RGB-D or LiDAR) and maintain communication channels for goal disclosure, introducing extra computation and potential privacy risks.
+While adaptable and fully decentralized, learning-based methods often require extensive training, may generalize poorly, and provide weaker guarantees on solution quality or completeness when relying solely on local information.
+Hybrid Methods: Hybrid planning–learning approaches aim to combine the strengths of both paradigms.
+FOLLOWER (Skrynnik et al.
+2024) uses congestion-aware $A^{*}$ planning to generate subgoals, followed by decentralized RL for replanning.
+DCC (Ma et al.
+2021b) and EPH (Tang et al.
+2024) augment local FOV sensing with selective inter-agent communication, while DHC (Ma et al.
+2021a) exchanges information with $K$ nearest neighbors to improve coordination.
+EPH further employs ensemble-based inference and switches to fast $A^{*}$ when no agents appear in FOV.
+More recent systems such as LNS2+RL (Wang et al.
+2025) combine LNS with MARL for early conflict resolution before refining solutions via search.
+Although effective, these hybrid designs often require complex integrations and specific information dependencies, and typically trade optimality for scalability.
+To enable a structured comparison across search-based, learning-based, and hybrid MAPF methods, we adopt a four-stage pipeline: S1 (Agent Planning), S2 (Collision Detection), S3 (Collision Avoidance Policy), and S4 (Agent Replanning).
+Within this view, CBS performs decentralized $A^{*}$ for S1/S4 and uses a centralized high-level solver for S2/S3, ensuring optimality.
+LNS centralizes all four stages, path generation, conflict detection, repair, and subset replanning, trading optimality for scalability.
+Learning-based approaches rely on local observations to decentralize S1–S4, though real deployments typically require onboard sensing (e.g., cameras, LiDAR) and inter-agent communication, increasing cost and computation.
+Hybrid methods occupy the middle ground: FOLLOWER uses global congestion maps with $A^{*}$ for S1 and decentralized RL for S2–S4, while LNS2+RL combines centralized LNS for S1–S3 with a hybrid RL/prioritized-planning strategy for S4.
+A detailed categorization is provided in Table 6 (Appendix Section A.4).
+Overall, existing MAPF paradigms face key trade-offs: centralized methods struggle with scalability and privacy; purely learning-based methods can weaken solution quality; and current hybrids still rely on substantial centralized coordination or costly sensing.
+The hybrid framework, IC–MAPF, introduced in Section 3 is designed to overcome these limitations by reducing inter-agent information sharing while preserving feasibility, thereby avoiding the communication and sensing overhead typical of learning-based systems.
 3 Methodology
-3.1 Approach Summary
-We propose a hybrid coordination framework that leverages decentralized RL-based path planning together with centralized collision detection and control.
-While conceptually similar to the high-level loop of Conflict-Based Search (CBS), our framework differs significantly in its resolution step: it does not build a constraint tree or perform backtracking, instead issuing targeted alerts that prompt heuristic-based replanning from the local RL agent.
-The proposed method’s four-stage MAPF pipeline is detailed below.
+3.1 Information-Centric MAPF
+Algorithm Category Information Used by Agent $i$
+CBS/ EECBS/ LaCAM/ LaCAM*/ LNS/ LNS2 Centralized (Search-Based) $\mathcal{I}_{i}^{\text{central}}(t)=(\mathcal{M},\,\mathcal{S}_{i}=(s_{i}^{t},g_{i}^{t},\pi_{i|t_{0}:t_{\max}}),\,\mathcal{O}_{j}=(s_{j}^{t},g_{j}^{t},\pi_{j|t_{0}:t_{\max}})_{j\neq i})$
+MRCDRL / AMPPO-PP Centralized (Learning-Based) $\mathcal{I}_{i}^{\text{central}}(t)=(\mathcal{M},\,\mathcal{S}_{i}=(s_{i}^{t},g_{i}^{t},\pi_{i|t_{0}:t_{\max}}),\,\mathcal{O}_{j}=(s_{j}^{t},g_{j}^{t},\pi_{j|t_{0}:t_{\max}})_{j\neq i})$
+PIBT / CS-PIBT / Distributed (Search-Based) $\mathcal{I}_{i}^{\text{distributed}}(t)=(\mathcal{M}_{i}^{k},\,\mathcal{S}_{i}=(s_{i}^{t},g_{i}^{t},\pi_{i|t_{s}:t_{r}}),\,\mathcal{O}_{j}=\{(s_{j}^{t},g_{j}^{t},\pi_{j|t_{s}:t_{r}})\}_{j\in\mathcal{N}_{i}(t)})$
+SACHA / PRIMAL / SCRIMP Distributed (Learning-Based) $\mathcal{I}_{i}^{\text{distributed}}(t)=(\mathcal{M}_{i}^{k},\,\mathcal{S}_{i}=(s_{i}^{t},g_{i}^{t},\pi_{i|t_{s}:t_{r}}),\,\mathcal{O}_{j}=\{(s_{j}^{t},g_{j}^{t},\pi_{j|t_{s}:t_{r}})\}_{j\in\mathcal{N}_{i}(t)})$
+FOLLOWER Decentralized (Learning-Based) $\mathcal{I}_{i}^{\text{decentral}}(t)=(\mathcal{M}_{i}^{k},\,\mathcal{S}_{i}=(s_{i}^{t},g_{i}^{t},\pi_{i|t_{s}:t_{r}}))$
+MA-CBS/ M*/ HiPP-MAPP/ DCC/ DHC/ EPH/ IC–MAPF Hybrid $\mathcal{I}_{i}^{\text{hybrid}}(t)=(\mathcal{M}_{i},\,\mathcal{S}_{i}=(s_{i}^{t},g_{i}^{t},\pi_{i|t_{s}:t_{r}}),\,\mathcal{O}_{j}=\{(s_{j}^{t},g_{j}^{t},\pi_{j|t_{s}:t_{r}})\}_{j\in\mathcal{N}_{i}(t)},\mathcal{E}_{ij})$
+Table 1: Categorization of MAPF algorithms by information availability, based on the unified model $\mathcal{I}_{i}(t)=(\mathcal{M}_{i},\mathcal{S}_{i},\mathcal{O}_{j})$.
+In an MAPF problem, each agent $i\in\{1,\dots,N\}$ relies on a set of information components to safely navigate from its start to its goal.
+This includes environmental information, represented by the entire map or a subset of the map $\mathcal{M}_{i}=(V,E,O)$ of vertices, edges, and static obstacles; self information, consisting of the agent’s current state $s_{i}^{t}\in V$, goal $g_{i}^{t}\in V$, and its augmented policy over a time interval $\{t_{s},...,t_{r}\}$ given by $\pi_{i|t_{s}:t_{r}}$, which can be compactly denoted by $\mathcal{S}_{i}=(s_{i}^{t},g_{i}^{t},\pi_{i|t_{s}:t_{r}})$; and some amount of other-agent information, specifically their current positions, goal positions and the immediate intended actions following their individual policies, $\mathcal{O}_{j}=(s_{j}^{t},g_{j}^{t},\pi_{j|t_{s}:t_{r}})_{j\in\mathcal{N}_{i}(t)}$, where $\mathcal{N}_{i}(t)$ denotes the set of agents in the neighborhood of agent $i$, required to avoid vertex and edge collisions.
+Collectively, the information accessible to agent $i$ at any timestep t, is written as $\mathcal{I}_{i}(t)=\bigl(\mathcal{M}_{i},\,\mathcal{S}_{i},\mathcal{O}_{j}\bigr).$ While this set captures the general information required for MAPF, different algorithmic paradigms make use of $\mathcal{I}_{i}(t)$ in distinct ways.
+In centralized algorithms, a global solver gathers the full map information and the full joint information $\{s_{i}^{t},\,g_{i}^{t}\}_{i=1}^{N}$ and computes conflict-free paths $\{\pi_{i}\}_{i=1}^{N}$; during execution, each agent requires only $\mathcal{I}_{i}^{\text{central}}(t)=(\mathcal{M}_{i},\mathcal{S}_{i},\mathcal{O}_{j})$ where $\mathcal{M}_{i}=\mathcal{M}$, and $t_{s}=t_{0}$ and $t_{r}=t_{max}$, where $\{t_{0},...,t_{max}\}$ denotes the entire time-interval required for the agent to traverse from its start to goal positions.
+Note that the central coordinator under this setting requires the information $O_{j},j\neq i$ of all the other agents, including the neighboring agents.
+In contrast, distributed algorithms require each agent to make decisions based solely on locally available information, leading to $\mathcal{I}_{i}^{\text{distributed}}(t)=\bigl(\mathcal{M}_{i}^{k},\mathcal{S}_{i},\mathcal{O}_{j}\bigr),$ where $\mathcal{M}_{i}$ may be a $k\times k$ subgrid of the map, often referred to as the FOV, $\mathcal{S}_{i}$ is the agent’s self information and $\mathcal{O}_{j}$, with $j\in\mathcal{N}_{i}$ is the neighboring agents’ information which is either sensed or exchanged.
+On the other hand, in decentralized algorithms, each agent only has access to the environment information and self-information, using which they independently plan out paths.
+So the information in these strategies can be denoted as $\mathcal{I}_{i}^{\text{decentral}}(t)=\bigl(\mathcal{M}_{i}^{k},\mathcal{S}_{i}\bigr).$
+Finally, hybrid algorithms combine global and local reasoning: a central module resolves large-scale couplings or computes partial plans, while agents use local information for real-time coordination, yielding $\mathcal{I}_{i}^{\text{hybrid}}(t)=\bigl(\mathcal{M}_{i},\,\mathcal{S}_{i},\mathcal{O}_{j},\mathcal{E}_{ij}\bigr),$ where additional information $\mathcal{O}_{j}$ related to other agents is shared to an agent only under the occurrence of specific events $\mathcal{E}_{ij}$, like collisions or deadlocks.
+Thus, although all MAPF settings require the same foundational information, they differ fundamentally in how much of it is shared, how widely it is distributed, and which components are used to compute safe and coordinated multi-agent behavior.
+To illustrate this, the information exchange across various categories of well known MAPF algorithms is presented in Table 1.
+3.2 Proposed Approach for IC–MAPF
+To address this problem of information-centric MAPF, we treat collision resolution as a mechanism for controlling when additional information about other agents needs to be shared.
+At a high level, a hybrid four-stage strategy can be built.
+In S1 (Decentralized Path Planning), each agent independently computes a plan from its start to its goal using information about the map and static obstacles.
+In S2 (Centralized Collision Detection), a coordinating module examines the collection of agent plans over time and identifies conflicts (vertex or edge conflicts).
+In S3 (Collision Control), a central control policy interprets each detected conflict and specifies a set of constraints to resolve the conflict.
+Finally, in S4 (Replanning), one of the agents involved in the conflict replans its path following the constraints from S3.
+By varying what constraints are set for replanning, we obtain a family of tiered replanning strategies (S4.1–S4.4) that trade off information usage against solution completeness.
+(S4.1) yield-based local coordination, where a single agent performs a short detour to a nearby parking location, waits, and then rejoins its original plan using only immediate (typically 1 to 3 hop distance) local occupancy information; (S4.2) static replanning, where a small segment of one agent’s plan is recomputed while treating a fixed set of conflict cells as forbidden and keeping all other plans unchanged; (S4.3) dynamic replanning, where a small segment is replanned while treating short prefixes of other agents’ plans as time-varying obstacles over a finite horizon; and (S4.4) local joint planning, where a tightly coupled subset of agents is replanned jointly over a bounded window while all remaining agents are treated as fixed reservations.
+Together, these tiers provide a controlled way to gradually increase the amount of information used for coordination only when simpler, lower inter-agent information-based repairs fail.
+We provide completeness and soundness properties of these individual strategies and the aggregated strategy we employed in this work in Appendix A.2.
+The proposed methodology of the four-stage MAPF pipeline is detailed below.
 S1: Decentralized Path Planning
@@
 $\rho_{k}=\pi_{\theta}(s_{k},g_{k})=(v^{k}_{0},\dots,v^{k}_{\tau_{k}}),$ (1)
-where $v^{k}_{0}=s_{k}$, $v^{k}_{\tau_{k}}=g_{k}$, and $\pi_{\theta}$ is a parameterized RL policy trained to minimize path length and local collision risk over a planning horizon $H$.
+where $v^{k}_{0}=s_{k}$, $v^{k}_{\tau_{k}}=g_{k}$, and $\pi_{\theta}$ is a parameterized RL policy trained to minimize path length and collision risk over a planning horizon $H$.
 No information about other agents is exchanged during this stage.
@@
 Once an alert is issued to an agent, the selected agent is prompted to perform a constrained replanning of its trajectory for a rollout window $\{t^{c_{k}}_{j-r},\dots,t^{c_{k}}_{j},\dots,\tau_{c_{k}}\}$, where $r$ is the rewind window by avoiding the collision set $\Delta_{c}$.
-S4: Decentralized Replanning
-Upon receiving a collision alert, agent $a_{c_{k}}$ truncates its path at $v^{c_{k}}_{t_{j-r}}$, denoted as
+S4: Replanning
+Upon receiving a collision alert for conflict time $t_{j}$ and rewind parameter $r$, the selected agent $a_{c_{k}}$ conceptually decomposes its original trajectory $\rho_{c_{k}}=(v^{c_{k}}_{0},\dots,v^{c_{k}}_{\tau_{c_{k}}})$ into a fixed prefix, a replanning window, and an untouched suffix.
+We denote the fixed prefix up to just before the rewind step as
 $\rho_{c_{k}|t_{j-r-1}}=(v^{c_{k}}_{0},\dots,v^{c_{k}}_{t_{j-r-1}}).$ (5)
-The agent then generates a new path segment $\rho_{c_{k}}^{\prime}$ from $v^{c_{k}}_{t_{j-r}}$ to its goal $g_{c_{k}}$ by applying its RL policy $\pi_{\theta}$.
-This replanning incorporates new constraints derived from the alert.
-For static obstacle avoidance (S4.1), the policy is typically derived by incorporating a fixed constraint, wherein the agent is supposed to avoid forbidden cells, as captured in Eq.
-6, where $\Delta_{c}$ represents the set of static cells identified in the alert, and a segment of the new path is generated as:
-$\rho_{c_{k}}^{\prime}=\pi_{\theta}(v^{c_{k}}_{t_{j-r}},g_{c_{k}}\mid v^{c_{k}}_{t_{i}}\notin\Delta_{c},i\geq j-r).$ (6)
-Alternatively, for dynamic obstacle avoidance (S4.2), the replanning must account for the predicted movements of other agents involved in the collision.
-In this case, the RL policy is conditioned on the sub-paths of the conflicting agents
-$\rho_{c_{k}}^{\prime}=\pi_{\theta}(v^{c_{k}}_{t_{j-r}},g_{c_{k}}\mid v^{c_{k}}_{t}\neq v^{c_{l}}_{t},v^{c_{l}}_{t}\in\rho_{c_{l}},$
-$\forall l\neq k,a_{c_{l}}\in A_{c},\forall t\in[t_{j-r},t_{j+r}]).$ (7)
+All S4 operators then construct a new path segment $\rho_{c_{k}}^{\prime}$ that replaces the original states on the interval $\{t_{j-r},...,t_{j+r}\}$.
+(S4.0) Yield-based local coordination.
+As a first step, the controller attempts a lightweight yield maneuver.
+If the conflict set $A_{c}$ contains an agent that is already parked at its start or has reached and is waiting at its goal at time $t_{j}$, we mark this agent as anchored and select it as the replanning agent $a_{c_{k}}$; otherwise, S3 selects the replanning agent $a_{c_{k}}\in A_{c}$ using a heuristic policy (e.g., farthest from goal).
+In both cases, the remaining agents in $A_{c}$ are treated as protected and keep their current plans.
+Starting from its rewind state $v^{c_{k}}_{t_{j-r}}$, $a_{c_{k}}$ then executes a short-horizon yield: it moves to a nearby parking cell, waits while the protected agent(s) traverse the conflict region, and then returns to a state from which it can resume progress toward $g_{c_{k}}$.
+The parking cell is chosen by a local search that ensures the resulting detour is free of vertex and edge conflicts with all other agents over the considered window.
+If no such parking cell exists, or if the yield trajectory does not eliminate the conflict when re-simulated by the controller, the plan is abandoned, and the system falls back to try the next strategy.
+For static obstacle avoidance (S4.1), the constraint set, $\Delta_{c}$ in Eq.
+6, is fixed by the alert and encodes the static cells that must be avoided.
+The “bounded” aspect comes from limiting how much of the trajectory is recomputed in time: we select a sub-initial point $v^{c_{k}}_{t_{j-r}}$ and a sub-goal (either the true goal $g_{c_{k}}$ or $v^{c_{k}}_{t_{j+r}}$), and only replan this suffix under the fixed constraint.
+The segment of the new path is generated as:
+$\rho_{c_{k}}^{\prime}=\pi_{\theta}(v^{c_{k}}_{t_{j-r}},v^{c_{k}}_{t_{j+r}}\mid v^{c_{k}}_{t_{i}}\notin\Delta_{c},\forall t\in\{t_{j-r},...,t_{j+r}\}).$ (6)
+If bounded static replanning still fails to resolve the conflict, the controller may share richer information and trigger dynamic obstacle avoidance (S4.2), where the replanning must account for the predicted movements of other agents involved in the collision.
+In this case, the RL policy is conditioned on the sub-paths of the conflicting agents as illustrated by the constraints in Eq.
+7
+$\rho_{c_{k}}^{\prime}=\pi_{\theta}(v^{c_{k}}_{t_{j-r}},v^{c_{k}}_{t_{j+r}}\mid v^{c_{k}}_{t}\neq v^{c_{l}}_{t},v^{c_{l}}_{t}\in\rho_{c_{l}},$
+$\forall l\neq k,a_{c_{l}}\in A_{c},\forall t\in\{t_{j-r},...,t_{j+r}\}).$ (7)
 The information guiding the replanning–whether it constitutes minimal details such as static obstacle constraints (Eq.
-6), or more detailed sub-path information about colliding agents’ paths treated as dynamic obstacles (Eq.
-7)–is integrated into the RL agent’s decision-making process (by modifying its state representation, more details in the Section 3.2) to guide it towards a conflict-free solution.
-The agent’s new complete trajectory $\rho^{new}_{c_{k}}$ is formed by concatenating the initial segment with the newly planned one, given by
-$\rho^{new}_{c_{k}}=\rho_{c_{k}|t_{j-r}}\|\rho_{c_{k}}^{\prime}.$ (8)
-This updated trajectory is submitted to the central controller.
-This iterative cycle of detection (S2), control (S3), and selective replanning (S4) continues until a globally conflict-free solution is achieved ($C(\rho,\tau)=\emptyset$).
-Our hybrid MAPF framework manages a precise flow of information, essential to support our claims of reduced information sharing and adaptive planning.
-The process begins with fully decentralized agent planning (S1), where each RL-driven agent uses only its local information before submitting its intended path $\rho_{k}$ to a central coordinator.
-This coordinator, leveraging its global observation of all submitted paths, performs collision detection (S2) to identify the set of all potential conflicts $C(\rho,\tau)$.
-Following detection, the central control module (S3) issues targeted alerts; it selects a specific agent $a_{c_{k}}$ and determines the rewind point $t_{j-r}$ from which the agent must replan, effectively issuing an alert $\mathcal{A}(c)$.
-S3 also governs the level of information to be shared for the subsequent decentralized replanning (S4).
-Initially, the alerted agent $a_{c_{k}}$ attempts to resolve the conflict using minimal information, typically by treating its own conflicting path segment $\Delta_{c}$ as a static obstacle (S4.1).
-If this localized, low-information approach proves insufficient, S3 then shares additional information with the agent–specifically, a subset of the path of other directly conflicting agents–for a more informed replan (S4.2).
-This tiered strategy—from purely local information in S1, to targeted alerts from S3, and then to progressively detailed yet still localized conflict information for S4—ensures that inter-agent information sharing is demand-driven and sparse.
-Agents communicate their revised plans $\rho^{new}_{c_{k}}$ back to the coordinator, and this iterative cycle of detection and selective, adaptive replanning continues until a globally conflict-free solution is achieved ($C(\rho,\tau)=\emptyset$).
-This methodology substantiates our framework’s ability to ensure feasibility while operating with significantly reduced information exchange.
+6) or more detailed sub-path information about colliding agents’ paths treated as dynamic obstacles (Eq.
+7)–is integrated into the RL agent’s decision-making process (by modifying its state representation, more details in Section 3.4) to guide it towards a conflict-free solution.
+In all cases, the agent’s new complete trajectory $\rho^{new}_{c_{k}}$ is formed by concatenating the initial segment with the newly planned one, given by
+$\rho^{new}_{c_{k}}=\rho_{c_{k}|t_{j-r-1}}\,\|\,\rho_{c_{k}}^{\prime}\,\|\,\rho_{c_{k}|t_{j+r+1}}.$ (8)
+Each updated trajectory is submitted to the central controller.
+The iterative cycle of detection (S2), control (S3), and decentralized replanning (S4) continues until no further conflicts can be resolved by single-agent strategies or a global termination condition is reached.
+In a small fraction of cases, even dynamic single-agent replanning leaves a persistent local conflict among a tightly coupled subset of agents.
+To address these situations, the controller invokes a local joint planner (S4.3) over a subset $A_{c}^{J}\subseteq A_{c}$ and a bounded time window around the conflict.
+Let $\{t_{j-r_{J}},...,t_{j+r_{J}}\}$ denote a short horizon centered at $t_{j}$.
+For each agent $a_{k}\in A_{c}^{J}$, we treat $v^{k}_{t_{j-r_{J}}}$ as its joint-planning start state and $v^{k}_{t_{j+r_{J}}}$ as a temporary sub-goal, and we solve a small-horizon problem in the joint configuration space of $A_{c}^{J}$ over this interval while treating all other agents as fixed reservations.
+The resulting joint segment $\{\rho_{k}^{\mathrm{joint}}\}_{a_{k}\in A_{c}^{J}}$ is spliced into the original trajectories between $t_{j-r_{J}}$ and $t_{j+r_{J}}$, analogous to Eq.
+8.
+If none of the above strategies can successfully remove a conflict, the corresponding agent is temporarily deferred: it is parked (typically at its start) and excluded from further replanning during the current phase.
+Once the non-deferred agents have reached a conflict-free configuration up to the current makespan, a second phase is initiated in which the deferred agents are reintroduced and planned using the same tiered strategies.
+This two-phase design specifically targets deadlock-like situations by decoupling stubborn local clusters and allowing them to be resolved after the rest of the system has stabilized.
+3.3 Execution Flow
+Our hybrid MAPF framework manages a precise flow of information, which is central to our goal of reducing inter–agent communication while preserving solution quality.
+The process begins with fully decentralized agent planning (S1), where each RL-driven agent uses only its local information to compute an intended path $\rho_{k}$ and submits this plan to a central coordinator.
+The coordinator, leveraging its global view of all submitted paths, performs collision detection (S2) to identify the set of potential conflicts $C(\rho,\tau)$.
+For each conflict, the control module (S3) issues a targeted alert $\mathcal{A}(c)$: it selects a replanning agent $a_{c_{k}}$, chooses a rewind window $\{t_{j-r},..,t_{j+r}\}$, and specifies which tiered strategy should be invoked.
+In S4, the selected agent first attempts to resolve the conflict using minimal information through a local yield maneuver (S4.0), which relies only on occupancy checks and a nearby parking cell.
+If the yield is not applicable or fails, S3 gradually increases the information budget: it provides a static constraint set derived from the conflict (S4.1), then, if needed, short sub-paths of conflicting agents are treated as dynamic obstacles over the same window (S4.2).
+When these single-agent repairs are insufficient, a local joint A* over a small group of agents and a bounded horizon is invoked (S4.3), with all remaining agents treated as fixed reservations.
+Agents that still cannot be repaired are temporarily deferred and handled in a second phase with relaxed attempt limits to break deadlocks.
+As shown in Appendix A.2, each tiered strategy in S4 is sound: whenever it returns a modified plan, the resulting joint execution remains collision-free over its replanned window.
+Moreover, under standard MAPF assumptions on finite grids with individually feasible start-goal pairs, the aggregated strategy with the deferral mechanism yields a complete repair procedure for our setting: every resolvable conflict is eventually eliminated, or else identified as belonging to an irreducible deadlock component among deferred agents.
+This tiered execution loop ensures that additional information is only shared on demand, allowing our framework to maintain feasibility while operating with significantly reduced information exchange.
 3.4 Policy Representation
 (S1 & S4) Decentralized Path Planning:
-In our proposed framework, we address the challenges of decentralized multi-agent pathfinding in dynamic environments by integrating collision awareness directly into the agent’s observation.
-Our approach leverages an RL-based decentralized planner that receives a multi-channel grid observation.
-Observation Space: Each agent’s observation is a tensor $\mathbf{s}\in\mathbb{R}^{H\times W\times 4}$, where $H$ and $W$ are the height and width of the grid.
-The four channels correspond to: (a) ObstacleMap - A binary map (1 if a static obstacle is present, 0 otherwise).
-(b) AgentMap - A binary map marking the agent’s current position.
-(c) GoalMap - A binary map indicating the goal position.
-(d) AlertMask - A dynamic collision alert map (initially all zeros, updated during execution).
-This alert mask is updated online by a centralized collision detection module that simulates timely collision alerts.
-In addition, the observation includes low-dimensional features comprising a unit vector pointing from the agent’s current position to the goal and the Euclidean distance.
-Action Space: The agent operates in a discrete action space, $\mathcal{A}=\{0,1,2,3,4\}$, corresponding to movements in the four cardinal directions (e.g., Up, Down, Left, Right, WAIT).
-Reward Structure Our reward function, similar to PRIMAL (Sartoretti et al.
-2019), promotes efficient, collision-free navigation and discourages unproductive behaviors.
-Agents receive +20 for reaching the goal.
-Penalties include: -3 for collisions with obstacles, -2 for timeout, -0.02 per time step taken (to encourage efficiency), and -0.1 for a ‘WAIT’ action, a penalty that encourages path progression over stationary waiting.
-An additional -0.05 penalty is applied if the agent is near a dynamic obstacle.
+Our framework employs an RL-based decentralized planner for dynamic replanning S4.2 that incorporates collision awareness directly into the agent’s observation, enabling effective path planning in dynamic multi-agent environments.
+Observation Space: Each agent observes a tensor $\mathbf{s}\in\mathbb{R}^{H\times W\times 4}$, where the channels encode: (a) a binary ObstacleMap for static obstacles, (b) an AgentMap for the agent’s current position, (c) a GoalMap denoting the goal location, and (d) an AlertMask, initially zero and updated online to reflect collision alerts generated by the centralized detection module.
+The observation is augmented with low-dimensional features consisting of a unit vector toward the goal and the Euclidean distance to the goal.
+Action Space: The agent operates in a discrete action space $\mathcal{A}=\{0,1,2,3,4\}$, corresponding to movements in the four cardinal directions and a WAIT action.
+Reward Structure: The reward function, following standard in learning-based approaches, encourages efficient, collision-free navigation.
+Agents receive a reward of +20 upon reaching the goal and are penalized for obstacle collisions (-3), timeouts (-2), each timestep (-0.02), and waiting actions (-0.1).
+An additional penalty of -0.05 is applied when the agent is near a dynamic obstacle.
 (S2) Rule-Based Collision Detection:
 The collision detection module (S2) functions as a deterministic, rule-based system.
-It takes the set of all current agent trajectories $\rho=\{\rho_{1},\dots,\rho_{n}\}$ and their makespans $\tau$ as input.
-For each timestep $t=\{0,1,...,T_{M}\}$ upto the maximum makespan ($T_{M}$), the module systematically scans for two primary types of conflicts: vertex conflicts (where $v^{k}_{t}=v^{\ell}_{t}$ for $k\neq\ell$) and edge conflicts (where $(v^{k}_{t},v^{k}_{t+1})=(v^{\ell}_{t+1},v^{\ell}_{t})$ for $k\neq\ell$).
-Upon detecting a conflict, characterized as $c=(t,v,A_{c})$ where $A_{c}$ is the set of agents involved, S2 reports this conflict $c$ to the S3 control module for resolution.
+It takes the set of all current agent trajectories $\rho=\{\rho_{1},\dots,\rho_{n}\}$ as input.
+For each timestep $t=\{0,1,...,T_{M}\}$, where $T_{M}$ is maximum makespan, the module systematically scans for vertex and edge conflicts, and reports these to the S3 control module for resolution.
 The detection process for all conflicts is computationally efficient, with a time complexity of $O\bigl(\sum_{k}|\rho_{k}|\bigr)$ per cycle, linear in the sum of all agent path lengths.
@@
 4 Experimental Setup
-In this section we first describe how we train the per‐agent navigation policy used in stage S1, then detail the maps, problem instances, and competing planners used to evaluate stages S2–S4, and define the performance metrics used to evaluate.
+In this section, we first describe how we train the per‐agent navigation policy used in stage S1, then benchmark dataset and planner details used to evaluate stages S2–S4, and define the performance metrics used to evaluate.
+We additionally provide details on the hardware experiments performed.
 Training Procedure
-We train a parametrized RL model $Q_{\theta}(s,a)$ using the DDQN algorithm(Van Hasselt, Guez, and Silver 2016), incorporating prioritized experience replay (PER) and $\varepsilon$‑greedy exploration, to enable an agent to navigate to a specified goal in the presence of both static and dynamic obstacles.
-Each dynamic obstacle possesses a hidden goal and follows precomputed trajectories, executing only valid moves; this allows for the simulation of online collision alerts during inference from the S3 stage.
-Training is performed for 30 00030\,000 episodes on an $11\times 11$ maze grid, with each episode capped at a maximum of $T_{\max}=50$ steps.
-The curriculum for environmental complexity is scheduled as follows: during the first 500 episodes, static‑obstacle density $\rho_{s}=0.10$ and dynamic‑obstacle count $n_{d}=0$; from episode 500 to 2999, $\rho_{s}=0.10$ and $n_{d}=1$; from episode 3000 to 5999, $\rho_{s}=0.20$ and $n_{d}=2$; and for episode 6000 onward, $\rho_{s}=0.30$ and $n_{d}=4$.
-Input observations are normalized to zero mean and unit variance before being fed to the network.
-Transitions $(s_{t},a_{t},r_{t},s_{t+1})$ are stored in a prioritized experience replay buffer of size $10^{6}$.
-At each learning step, a mini‑batch of $128$ transitions is sampled.
-The agent selects actions using an $\varepsilon$‑greedy policy, where $\varepsilon$ decays from $\varepsilon_{0}=1.0$ to $0.01$ over the course of training according to $\varepsilon_{t+1}=\max(0.01,0.999\,\varepsilon_{t})$.
-Crucially, during both exploration and exploitation (e.g., when calculating the target Q-value), only actions $a_{t}$ deemed valid by an action mask $m_{t}\in\{0,1\}^{|\mathcal{A}|}$ are considered for efficient exploration (Damani et al.
-2021b).
-The Q-network parameters $\theta$ are updated via gradient descent using the Adam optimizer with a learning rate $\alpha=3\times 10^{-4}$.
-The discount factor is $\gamma=0.97$.
-A separate target network, with parameters $\theta^{-}$, provides stable targets and is updated by copying $\theta$ every 300 steps.
-Updates minimize the prioritized, importance‑weighted Bellman error.
-The target value $y_{t}$ and TD-error $\delta_{t}$ are computed as: $y_{t}=r_{t}+\gamma\,Q_{\theta^{-}}\bigl(s_{t+1},\arg\max_{a^{\prime}}Q_{\theta}(s_{t+1},a^{\prime})\text{ s.t.
-}m_{t}(a^{\prime})=1\bigr),$ $\delta_{t}=y_{t}-Q_{\theta}(s_{t},a_{t}),$ The loss function is then calculated as $L(\theta)\;=\;\mathbb{E}_{i\sim p(i)}\bigl[w_{i}\,\delta_{i}^{2}\bigr],$ where the sampling probability $p(i)\propto|\delta_{i}|^{\alpha}$ ($\alpha$ is a hyperparameter for PER) and $w_{i}$ are importance‑sampling weights that correct for the bias introduced by prioritized sampling.
-In addition to the DDQN model, a Proximal Policy Optimization (PPO) based model was also trained; details regarding its architecture (including the one used for DDQN model) and training procedure, along with comparative performance, are provided in the Appendix Section A.3.
-The DDQN model was selected due to its superior performance over the PPO model - results are presented in Table 3.
-Evaluation Scenarios and Baselines
-We evaluate our framework’s performance and scalability across three distinct grid configurations, representing two primary map types: maze and warehouse.
-All evaluations use ten randomly generated problem instances per configuration.
-For the maze map type, we test on an $11\times 11$ grid with approximately 45% static obstacles (5 to 20 agents) and a larger $21\times 21$ grid with approximately 35% static obstacles (32, 64, and 96 agents).
-For the warehouse map type, we utilize a $25\times 25$ grid layout with approximately 24% static obstacles as shelves (32, 64, and 96 agents).
-Our proposed hybrid framework is evaluated through two variants, Alert-BFS and Alert-A*, which are named based on their RL policy rollout strategy (Breadth-First Search or weighted A-star search, respectively).
-The impact of our specific replanning strategies is further analyzed in an ablation study in Appendix Table 5.
-These are compared against search-based planners, Conflict-Based Search (CBS) (Sharon et al.
-2015) and Improved CBS (ICBS) (Boyarski et al.
-2015), and leading learning-based online planners PRIMAL (Damani et al.
-2021a) and SCRIMP (Wang et al.
-2023) with the MAPF formulation of disappear at goal.
-While our framework variants operate within our proposed four-stage pipeline (S1–S4), the baseline methods are executed as standalone algorithms.
-Per-instance time limits are 50 seconds for $11\times 11$ mazes, 30 minutes for $21\times 21$ mazes, and 1 hour for $25\times 25$ warehouse scenarios.
-For online planners, the step limits are 256 for the $11\times 11$ maze, and 512 for both the $21\times 21$ maze and the $25\times 25$ warehouse, as detailed in Table 1.
-Additional results on various other map configurations are provided in the Appendix Section A.4.
+We train a parametrized Q-network $Q_{\theta}(s,a)$ using Double DQN (Van Hasselt et al.
+2016) with prioritized experience replay (PER) and $\varepsilon$-greedy exploration.
+The policy is trained to reach a specified goal while avoiding both static and dynamic obstacles.
+Dynamic obstacles follow valid precomputed trajectories with hidden goals, enabling the simulation of online collision alerts during inference from the S3 stage.
+Training is conducted for 30,00030{,}000 episodes on an $11\times 11$ maze, with each episode capped at $T_{\max}=50$ steps.
+Environmental difficulty is increased through a curriculum: the first 500 episodes use static obstacle density $\rho_{s}=0.10$ with no dynamic obstacles; episodes 500–2999 use $\rho_{s}=0.10$ with one dynamic obstacle; episodes 3000–5999 use $\rho_{s}=0.20$ with two dynamic obstacles; and all remaining episodes use $\rho_{s}=0.30$ with four dynamic obstacles.
+Observations are normalized to zero mean and unit variance and stored as transitions $(s_{t},a_{t},r_{t},s_{t+1})$ in a PER buffer of size $10^{6}$.
+Mini-batches of size 128 are sampled for learning.
+Action selection follows an $\varepsilon$-greedy policy, with $\varepsilon$ decayed from $\varepsilon_{0}=1.0$ to $0.01$ according to $\varepsilon_{t+1}=\max(0.01,0.999\,\varepsilon_{t})$.
+At all times, action selection and target computation are constrained by a validity mask $m_{t}\in\{0,1\}^{|\mathcal{A}|}$, ensuring that only feasible actions are considered (Damani et al.
+2021).
+Network parameters are optimized using Adam with learning rate $\alpha=3\times 10^{-4}$ and discount factor $\gamma=0.97$.
+A separate target network with parameters $\theta^{-}$ is updated every 300 steps.
+The Double DQN target is computed as $y_{t}=r_{t}+\gamma\,Q_{\theta^{-}}(s_{t+1},\arg\max_{a^{\prime}}Q_{\theta}(s_{t+1},a^{\prime})\ \text{s.t.}\ m_{t}(a^{\prime})=1)$, and the temporal-difference error as $\delta_{t}=y_{t}-Q_{\theta}(s_{t},a_{t})$.
+Training minimizes the PER-weighted Bellman loss $L(\theta)=\mathbb{E}_{i\sim p(i)}[w_{i}\,\delta_{i}^{2}]$, where $p(i)\propto|\delta_{i}|^{\alpha}$ and $w_{i}$ are importance-sampling weights.
+Simulation: Evaluation Scenarios and Baselines
+We evaluate IC–MAPF on standard benchmark maps from the Moving AI Lab dataset (Stern et al.
+2019b), including random grids ($32\times 32$, $64\times 64$ with 20% obstacles) and structured maps (den312d and warehouse).
+Agent counts range from 10–128 on random maps and 8–128 on structured maps.
+For each configuration, we generate 20 random instances and report aggregated results in Table 3.
+We compare IC–MAPF against representative search- and learning-based baselines: CBS, DCC, SCRIMP, and EPH.
+IC–MAPF is given a 10-minute limit per instance (Python implementation), CBS is capped at 120 seconds, and learning-based solvers are run with step limits (128 for random maps, 256 for den312d/warehouse) and a 5-minute wall-clock bound.
+Section 5 presents comparative results highlighting IC–MAPF’s scalability, solution quality, and communication efficiency.
 Performance Criteria
@@
 For successful instances, Makespan (MS) measures the time until the last agent reaches its goal.
-The number of Collisions (CO) measures the total count of path disagreements detected and successfully resolved by the framework to produce the final collision-free solution.
-A non-zero CO in a successful run reflects the planner’s total conflict resolution workload.
-Finally, total Time (T) measures the computational cost until a solution is found or the time limit is exceeded.
-Learning-based methods are evaluated by the average number of steps taken in the environment.
+To quantify information sharing, we define an abstract Information Unit (IU) as the data required to represent one agent’s state (e.g., its coordinates) at a single timestep.
+For learning-based methods, IU is estimated by counting the number of other agents observed within an agent’s field of view at each timestep, aggregated over all agents and time.
+This unified abstraction enables direct comparison of information load across algorithms.
+For IC–MAPF, each cell-level constraint issued in S3 for S4 replanning counts as one IU.
+For learning-based methods, IU reflects the other-agent information visible within each agent’s FOV, and we additionally count any inter-agent communication (as in DCC and EPH).
+We omit IU for CBS because it uses full global information.
+Hardware Experiments: TurtleBot4
+We evaluate our approach on five TurtleBot4 robots operating in a $6{\times}6$ indoor grid.
+Five problem instances are executed using a ROS 2 discovery server to ensure stable communication.
+Robots interface with the central controller only through a shared synchronization channel, with all other ROS topics kept isolated and no onboard perception used.
+At the start, each robot resets its pose via the ROS 2 reset_pos service.
+Execution proceeds in strict lockstep: at each step, a robot (i) performs the required in-place rotation and acknowledges completion, and then (ii) executes a fixed $0.45$ m forward motion and acknowledges again.
+The controller advances the system only when all robots have completed both actions, ensuring fully synchronized multi-robot execution.
+Details on test maps, setup images, and robot execution metrics are provided in Appendix Figures 45 and Tables 45.
+Search-based Hybrid Learning-based Solvers
+Map $m$ CBS IC–MAPF SCRIMP DCC EPH
+[…]
+[Uncaptioned image] random $\times32$-20 10-30 100 47.6 100 54.0 29.6 100 50.3 19x 100 55.5 11x 98 57.4 9.0x
+40-60 65 85.3 100 98.9 161.5 95 62.1 12x 85 82.1 23x 97 77.4 22x
+70-90 0 - 98 73.4 1824.2 77 83.8 2.6x 50 111.3 10x 53 114.5 12x
+100 0 - 85 76.5 3553.8 60 97.2 2.0x 40 117.5 11x 50 121.0 12x
+128 0 - 0 - - 70 96.6 - 0 128.0 - 5 127.5 -
+[Uncaptioned image] random $\times64$-20 10-30 85 106.3 100 113.9 10.2 23 121.9 137x 97 105.9 7.9x 98 104.1 6.5x
+40-60 42 130.0 100 125.6 88.5 0 128.0 - 83 116.8 9.5x 98 114.0 7.3x
+70-90 0 - 98 129.1 213.5 0 128.0 - 43 125.0 14x 85 121.1 11x
+100 0 - 90 130.1 462.1 0 128.0 - 25 127.2 13x 45 125.3 11x
+128 0 - 5 132.0 607.6 0 128.0 - 0 128.0 - 15 127.7 20x
+[…]
+16 85 128.7 100 124.0 21.8 0 256.0 - 100 130.7 8.8x 100 122.2 5.9x
+32 20 148.0 100 136.7 105.2 0 256.0 - 85 163.5 11x 100 127.0 6.0x
+64 0 - 100 153.1 759.9 0 256.0 - 90 190.7 11x 100 146.1 8.8x
+128 0 - 0 - - 0 256.0 - 5 254.8 - 90 207.7 -
+[Uncaptioned image] warehouse $1×\times6$3 8 100 175.7 100 192.8 0.8 0 256.0 - 100 191.8 5.8x 100 191.6 2.3x
+16 100 196.2 100 215.1 1.0 0 256.0 - 100 213.3 16x 100 212.9 11x
+32 15 221.3 100 220.3 10.7 0 256.0 - 100 219.2 8.3x 100 218.1 5.4x
+64 0 - 100 227.6 23.6 0 256.0 - 100 227.6 15x 100 227.8 11x
+128 0 - 100 234.7 139.1 0 256.0 - 95 236.8 14x 90 239.1 11x
+[…]
 5 Results and Discussion
 Here we present an empirical study, evaluating our central hypothesis: a MAPF framework with strategically reduced information sharing can achieve robust performance.
-We address our RQs concerning minimal information needs and comparative performance against approaches with different information paradigms using the experimental results presented in Table 1.
+We address our RQs concerning minimal information needs, and Table 3 summarizes the comparative performance of IC–MAPF against representative search-based (CBS) and learning-based (DCC, SCRIMP, EPH) MAPF solvers across random grids and structured maps.
+Additional experimental results are presented in Appendix Section A.1 and detailed results with runtime in Figure 3.
 (RQ1): Given the observation constraints of a decentralized setup, can an effective MAPF algorithm be created with one agent knowing nothing about other agents?
@@
 Ans.
-No—if agents have zero knowledge of one another, MAPF fails, losing completeness guarantees; but solutions are possible with targeted, minimal information sharing.
-Table 5 in the appendix presents comparative results for various inter-agent information sharing settings.
-Now, we turn our attention to the strategically limited minimal information sharing case.
-In our framework, strategically informed alerts convey either (S4.1) static constraints (the conflicting cell) or (S4.2) slightly richer, yet still localized, dynamic constraints (the sub-path of only the directly conflicting agent(s) for a short horizon).
-Table 1 shows performance comparison of considered methodologies across 11x11 maze (a), 21x21 maze (b), and $25\times 25$ (c) warehouse maps, respectively.
-As shown in Table 1, the performance of our framework, where agents initially plan with only local information and receive minimal, targeted alerts from a central coordinator upon conflict, achieves high Success Rates (SR).
-Specifically, on the $11\times 11$ and $21\times 21$ maze grids, SRs are high (90%-100%).
-On the more structured $25\times 25$ warehouse map (Table 1(c)), Alert-BFS and Alert-A* also achieve 100% SR for 32 and 64 agents, with Alert-A* maintaining a 60% SR in the dense agent scenario (96 agents), demonstrating effectiveness across different environment types with this minimal information sharing strategy.
-This success with minimal inter-agent information contrasts sharply with alternatives.
-Centralized planners (CBS, ICBS) demand complete global knowledge of all agents’ paths, a requirement that, while enabling optimality, leads to their low SR due to the combinatorial explosion in conflict resolution within practical time limits.
-In contrast, distributed learning frameworks (PRIMAL & SCRIMP) often operate under the assumption that each planning agent possesses full observation, of other agents and their goals, in a local field of view.
-While this localized observation can facilitate decentralized decision-making, it implies a continuous and potentially substantial information gathering (e.g., via onboard cameras or LiDAR sensors) and processing requirement for each agent.
-The success of our method across maze and warehouse environments underscores that this on-demand information is sufficient for high SR, thereby avoiding the continuous, dense local information exchange.
-Our quantitative analysis, detailed in the Appendix A.1, shows that this on-demand alert mechanism reduces the total information load by $\sim$93% compared to a continuous-observation distributed paradigm in this scenario.
-Table 1: Comparison of performance metrics for different methods on an $11\times 11$ grid (5–20 agents, $\sim$45% obstacles), a $21\times 21$ grid (32, 64, and 96 agents, $\sim$35% obstacles), and a $25\times 25$ warehouse grid (32, 64, 96 agents, $\sim$24% obstacles).
-Results are averaged over 10 problem instances per configuration.
-Key metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T), with Time averaged over all trials.
-For learning-based methods, the average steps taken while solving the problem instances are presented in brackets along with average wall-clock time taken.
-(a) 11x11 maze with 50 sec. time limit per problem instance ($256$ steps for learning-based methods)
-SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (s) ($\downarrow$)
-Methods 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20
-Alert-BFS 100.00% 100.00% 98.00% 17.8 23.6 27.4 9 53 191 1.1 4.2 11.8
-Alert-A* 100.00% 96.00% 82.00% 16.1 20.7 24.0 10 59 173 3.3 12.0 25.8
-CBS 85.00% 34.00% 2.00% 13.8 14.3 17.0 559 8785 3895 8.3 30.7 35.3
-ICBS 85.00% 26.00% 10.00% 14.0 14.8 12.8 885 3486 1466 5.9 25.1 28.6
-PRIMAL 61.67% 48.00% 28.00% 71.7 89.9 104.8 - - - 1.4 (142) 3.4 (176) 6.1 (214)
-SCRIMP 95.00% 70.00% 60.00% 19.3 28.6 38.8 1 2 8 0.4 (31) 1.6 (97) 3.9 (124)
-(b) 21x21 maze with 30 min. time limit per problem instance ($512$ steps for learning-based methods)
-SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
-Methods 32 64 96 32 64 96 32 64 96 32 64 96
-Alert-BFS 100.00% 90.00% 0.00% 86.6 131.4 - 528 18909 - 2.7 23.9 30.0
-Alert-A* 100.00% 80.00% 0.00% 41.4 72.0 - 457 11379 - 3.3 23.5 30.0
-CBS 0.00% 0.00% 0.00% - - - - - - 30.0 30.0 30.0
-ICBS 0.00% 0.00% 0.00% - - - - - - 30.0 30.0 30.0
-PRIMAL 0.00% 0.00% 0.00% - - - - - - 0.8 (512) 2.1 (512) 1.8 (512)
-SCRIMP 100.00% 20.00% 10.00% 65.9 135.0 94.0 15 220 357 0.1 (66) 1.0 (437) 2.4 (470)
-(c) 25x25 warehouse with 1 hour time limit per problem instance ($512$ steps for learning-based methods)
-SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
-Methods 32 64 96 32 64 96 32 64 96 32 64 96
-Alert-BFS 100.00% 100.00% 10.00% 94.3 125.9 96.0 237 6379 26553 5.2 28.0 59.5
-Alert-A* 100.00% 100.00% 60.00% 40.5 43.7 50.8 211 3129 16277 4.7 19.2 51.8
-CBS 0.00% 0.00% 0.00% - - - - - - 60.0 60.0 60.0
-ICBS 10.00% 0.00% 0.00% 38.0 - - 123107 - - 57.9 60.0 60.0
-PRIMAL 0.00% 0.00% 0.00% - - - - - - 1.0 (512) 3.1 (512) 6.6 (512)
-SCRIMP 100.00% 100.00% 90.00% 60.2 49.0 101.7 2 14 161 0.1 (60) 0.3 (49) 1.1 (143)
+No–A purely decentralized setting in which agents have no information about others is insufficient once interactions occur, as conflicts cannot be consistently resolved.
+However, our results show that effective MAPF does not require continuous observation or global plan sharing.
+Instead, solutions can be achieved with minimal, event-triggered information sharing consistent with our hybrid information model $\mathcal{I}_{i}^{\text{hybrid}}(t)$, where additional other-agent information is revealed only upon conflict events.
+Across unstructured random grids, IC–MAPF achieves high success rates while using orders of magnitude less information than learning-based baselines.
+On the random-$\times32$-20 map, IC–MAPF maintains $100\%$ SR up to 80 agents, with cumulative IU values ranging from 3.6 to 214.2.
+Under the same settings, SCRIMP requires 262–2823 IU and EPH requires 19.7–6155 IU, corresponding to approximately $3\times$ to over $30\times$ higher information usage, while DCC requires 26.1–6337.2 IU (about $7\times$$30\times$ IC–MAPF over the same range).
+On the larger random-$\times64$-20 map, IC–MAPF sustains $100\%$ SR up to 80 agents with IU below 212, whereas SCRIMP fails beyond 30 agents and exceeds 6000 IU, and EPH exhibits reduced SR while using $6\times$ to $13\times$ more information than IC–MAPF; DCC similarly achieves moderate SR but uses 10.6–2947.4 IU, i.e., roughly $7\times$$14\times$ more information.
+Similarly, on structured benchmarks, den312d, IC–MAPF achieves $100\%$ SR for 8–64 agents with IU between 3.6 and 759.9, while EPH requires up to 6722.3 IU to obtain comparable SR, and SCRIMP fails for 16 agents and above; DCC lies in between, using 22.6–8180.1 IU (approximately $6\times$$11\times$ IC–MAPF for 8–64 agents).
+On the warehouse map, IC–MAPF solves all tested instances (8–128 agents) with IU between 0.8 and 139.1, whereas EPH requires 1.7–1502.4 IU to achieve similar SR, and DCC uses 4.3–2016.5 IU (about $5\times$$16\times$ IC–MAPF), while SCRIMP achieves $0\%$ SR despite using several thousand IU.
+These results identify targeted, conflict-triggered sharing of short agent-path fragments over bounded horizons as the minimal information required for effective MAPF, enabling reduced total information load of up to one to two orders of magnitude.
 (RQ2): How does the proposed hybrid method compare to leading alternative search- and learning-based approaches in terms of performance, solution quality, and scalability?
 Ans.
-Our hybrid framework exhibits a compelling balance of performance and scalability with distinct trade-offs against established search-based and learning-based methods, as detailed in Table 1.
-Our Python-based framework demonstrates higher SR and scalability compared to search-based optimal solvers.
-CBS/ICBS (C++ based) shows constantly lower SR with increasing environment complexity, failing to solve any instances on the $21\times 21$ maze (for 32+ agents) and the $25\times 25$ warehouse map.
-While CBS/ICBS achieves optimal MS when they solve (Table 1(a)), their reported Collision (CO) counts (Table 1(a)) for such cases are very high; our methods report significantly fewer collisions when they succeed on these smaller maps.
-PRIMAL exhibits low SR across all tested configurations, often hitting step limits without success.
-SCRIMP is a strong learning-based performer with fast inference times and good MS.
-On the $11\times 11$ and $21\times 21$ mazes, its SR is comparable or better than our methods in some low-density cases, but can decline in denser scenarios ($21\times 21$, 64 agents: SCRIMP 20% SR vs.
-Alert-BFS 90% SR).
-On the $25\times 25$ warehouse map, SCRIMP maintains 100% SR for 32 and 64 agents, and achieves 90% SR for 96 agents, outperforming Alert-BFS (10% SR) and Alert-A* (60% SR) at this highest agent count in terms of success.
-However, SCRIMP also demonstrates significantly lower CO values than our methods across all successful warehouse instances attributing to its efficient multi-agent training.
-A critical factor for online planners is the offline training investment.
-For instance, PRIMAL’s extensive training on varied map sizes involved approximately 20 days on significant computational resources (Damani et al.
-2021a).
-SCRIMP, with its sophisticated transformer-based architecture and PPO training, also implies a substantial training regimen across diverse configurations to achieve its reported performance (Wang et al.
-2023).
-In contrast, our RL planning model, trained only on $11\times 11$ maze grids, was developed with a comparatively modest regimen: training was conducted on a Mac Mini equipped with an Apple M4 Pro chip (14-core CPU, 20-core GPU, 16-core Neural Engine) and 48GB unified memory, running for approximately 6 to 8 hours.
-Its subsequent effective performance when tested on different map types (map and warehouse) and larger grid sizes (up to $25\times 25$) indicates strong generalization from this simpler training setup.
+Table 3 comparison highlights three key dimensions: success rate (SR), makespan (MS), and total information load (IU).
+Search-based solvers such as CBS rely on full joint information $\mathcal{I}_{i}^{\text{central}}(t)$ and attempt to resolve all couplings through centralized reasoning.
+While optimal when successful, CBS fails to scale under the imposed time limits and does not solve most medium-to-large instances reported in the table.
+This illustrates the practical limitations of centralized planners as problem size grows.
+Learning-based distributed methods, including DCC, SCRIMP, and EPH, operate under $\mathcal{I}_{i}^{\text{distributed}}(t)$, where each agent continuously accesses local observations of other agents within a field of view or via explicit message passing.
+These methods show strong performance in some dense scenarios (e.g., EPH on warehouse), but incur substantially higher information loads.
+In Table 3, their IU values are reported as multipliers relative to IC–MAPF, often ranging from one to two orders of magnitude higher due to continuous perception or communication.
+In contrast, IC–MAPF adopts a hybrid execution strategy.
+This allows IC–MAPF to scale favorably across both random and structured maps, achieving high SR with competitive makespans while keeping information exchange sparse and bounded.
+The gradual escalation of information, from static constraints to short dynamic sub-paths, and finally to bounded joint replanning only when necessary, ensures that increased coordination cost is incurred only in genuinely hard cases.
+Overall, IC–MAPF occupies a middle ground between centralized optimal planners and fully distributed learning-based planners.
+These results demonstrate that carefully controlled, event-driven information sharing can provide a favorable trade-off between performance and communication complexity in MAPF.
 6 Conclusions
-Our hybrid framework combines decentralized RL-based planning with only minimal, targeted coordinated alerts–static conflict cell flags or brief conflict tracks–to maintain better SR across various challenging scenarios.
-While it incurs more collisions (CO) and sometimes longer makespan (MS) compared to highly specialized methods like SCRIMP, and its SR declines in extremely dense scenarios, it often delivers better overall feasibility and scalability, in scenarios with increasing agent counts, than purely search‑based methods and PRIMAL by strategically sharing minimum inter-agent information.
-Effective multi-agent coordination with reduced information exchange, as explored in this research, will lead to privacy-aware automated systems, enabling broader use in complex privacy-preserving applications like autonomous driving.
-Despite these promising results, our work has limitations.
-The current framework, optimized for disappear at target setting, may face challenges in stay at target behavior.
-Future work will investigate the minimal information necessary to effectively manage these stay at target scenarios, aiming to extend the framework’s applicability without compromising its core principle of sparse information exchange.
-Additionally, an ablation study of richer AlertMask encodings, like dynamic occupancy maps, could improve replanning efficiency.
+In this paper, we introduced the problem of Information-centric MAPF and proposed a hybrid framework, IC–MAPF, that significantly reduces the inter-agent information required for successful multi-agent coordination.
+We further defined a metric, Information Units (IU), to quantify the amount of other-agent information consumed during planning.
+Across a range of challenging MAPF benchmarks, IC–MAPF achieves high success rates while reducing information usage by $2\times$ to $23\times$ compared to state-of-the-art communication-efficient baselines.
+The broader implications of this work extend to privacy-aware autonomous systems, where reducing inter-agent information exchange is essential.
+By showing that multi-agent coordination is achievable with sparse, event-triggered information, our results support the development of scalable, bandwidth-efficient, and privacy-preserving multi-robot systems.
+Despite these promising results, our work has limitations: IC–MAPF may face challenges in extremely dense settings where conflicts are frequent, and its tiered strategy is currently hand-designed.
+As future work, we aim to explore learning-based mechanisms that automatically adapt the tier-selection policy and investigate reinforcement learning under explicit information constraints, enabling agents to optimize performance while minimizing IU usage.
 Appendix A Appendix
 Appendix Contents
-A.1.
-A Quantitative Model of Information Load ........................................................................................................................................................................A.1
-A.2. Literature review ........................................................................................................................................................................A.2
-    Literature Categorization Table 2 ........................................................................................................................................................................2
-A.3. Training Methods ........................................................................................................................................................................A.3
+A.1. Additional Results ........................................................................................................................................................................A.1
+  Detailed Performance Table – Table 3 ........................................................................................................................................................................3
+  Strategy Usage Summary – Figure 2 ........................................................................................................................................................................2
+  Detailed Performance Plots – Figure 3 ........................................................................................................................................................................3
+A.2.
+Properties of Tiered Replanning Strategies ........................................................................................................................................................................A.2
 A.3.
-PPO Training Procedure ........................................................................................................................................................................A.3
-  A.3.
-Neural Network Architecture ........................................................................................................................................................................A.3
-  A.3. Model Comparison ........................................................................................................................................................................A.3
+Hardware Results: TurtleBot4 ........................................................................................................................................................................A.3
+  Robot Grid Configurations – Figure 4 ........................................................................................................................................................................4
+  Physical Setup Image – Figure 5 ........................................................................................................................................................................5
+  Communication Statistics – Table 4 ........................................................................................................................................................................4
+  Execution Quality Statistics – Table 5 ........................................................................................................................................................................5
+A.4. Literature review ........................................................................................................................................................................A.4
+    Literature Categorization Table 6 ........................................................................................................................................................................6
+A.5. Training Methods ........................................................................................................................................................................A.5
+  A.5.
+PPO Training Procedure ........................................................................................................................................................................A.5
+  A.5.
+Neural Network Architecture ........................................................................................................................................................................A.5
+  A.5. Model Comparison ........................................................................................................................................................................A.5
     DDQN training performance - Figure 6 ........................................................................................................................................................................
     PPO training performance - Figure 7 ........................................................................................................................................................................
-  A.3. Model Comparison ........................................................................................................................................................................A.3
-    Table 3 ........................................................................................................................................................................
-A.4. Additional Experiments ........................................................................................................................................................................A.4
-  A.4.
-Agent Selection Policy Comparison ........................................................................................................................................................................A.4
-    Table 4 ........................................................................................................................................................................4
-  A.4.
-Inter-agent Information Sharing Comparison ........................................................................................................................................................................A.4
-    Table 5 ........................................................................................................................................................................5
-  A.4.
-Benchmark Dataset Evaluation ........................................................................................................................................................................A.4
-    Table 6 ........................................................................................................................................................................6
-A.1 A Quantitative Model of Information Load
-To quantitatively support our claim of reduced information sharing, we developed a simplified model to estimate the ”Total Information Load” of different coordination paradigms.
-This model is based on a simple Information Unit (IU), defined as the data required to represent one agent’s state (e.g., its coordinates) at a single timestep.
-Model Assumptions
-Our calculation is based on the 11x11 maze scenario with 20 agents, using metrics from the successful Alert-BFS runs in Table 1(a).
-  1. 1.
-    Scenario Data: We use a 20-agent instance (N=20) solved with a makespan (T) of 27 timesteps and requiring the resolution of 191 conflicts (C).
-  2. 2.
-    Average Path Length (L): We assume an average initial path length of 25 steps per agent.
-  3. 3.
-    Agent Density (D): For a distributed method like PRIMAL with a 10x10 FoV on an 11x11 grid, each agent has near-global vision.
-Therefore, each agent continuously observes all other N-1 agents.
-We set the average density (D) to 19 neighbors.
-  4. 4.
-    Alert Size ($I_{\text{alert}}$): We assume each conflict alert sent by our coordinator is a minimal, static alert costing 1 IU.
-Comparative Calculation
-Distributed Method (e.g., PRIMAL)
-The information load is the continuous sensing and processing of all visible neighbors by every agent at every timestep. The formula is $\text{Info}_{\text{Distributed}}=N\times T\times D$.
-$\text{Info}_{\text{Dist.}}$ $=20\times 27\times 19$
-$=\textbf{10,260 IU}$
-Our Hybrid Method (Alert-X)
-The load is the one-time path sharing in addition to all subsequent targeted alerts. The formula is $\text{Info}_{\text{Alert-X}}=(N\times L)+(C\times I_{\text{alert}})$.
-$\text{Info}_{\text{Alert-X}}$ $=(20\times 25)+(191\times 1)$
-$=500+191$
-$=\textbf{691 IU}$
+A.1 Additional Results
+Tables 3, Figures 2, and Figure 3 provide expanded performance analyses complementing the compact results presented in the main paper.
+Table 3 reports the full per-configuration statistics for all solvers across four benchmark maps, including success rate, makespan, and absolute values for Information Units (IU).
+These detailed values highlight the consistency of IC–MAPF across increasing agent densities and show the scale at which learning-based methods begin to degrade or saturate, especially in larger and more structured environments.
+Figure 2 summarizes the usage frequency of our tiered replanning strategies S4.0 to S4.3 (Yield, Static, Dynamic, and Joint) across all problem instances.
+This visualization illustrates how IC–MAPF automatically escalates strategy complexity only when needed, typically in high-density or tightly coupled regions, while relying predominantly on low-information repairs in easier configurations.
+Figure 3 presents aggregate trends for success rate, makespan, IU consumption, and average runtime.
+The shaded regions indicate standard deviation across 20 problem instances, providing insight into solver stability and variance across different metrics.
+Compared to CBS and learning-based planners (DCC, SCRIMP, EPH), IC–MAPF achieves consistently high success rates while maintaining substantially lower information usage and competitive makespan behavior.
+Together, these additional results reinforce the main claims of the paper: IC–MAPF offers a favorable balance between scalability, robustness, and dramatically reduced inter-agent information requirements.
+Search-based Hybrid Learning-based Solvers
+Map $m$ CBS IC–MAPF SCRIMP DCC EPH
+[…]
+[Uncaptioned image] random $\times32$-20 10 100 46.8 100 50.5 3.6 100 48.0 262.2 100 49.1 26.1 100 47.6 19.7
+20 100 47.6 100 56.0 18.8 100 50.5 565.2 100 58.0 211.2 95 62.5 204.2
+30 100 48.4 100 55.5 66.3 100 52.4 884.8 100 59.5 696.2 100 62.0 576.8
+40 90 80.0 100 107.0 48.8 95 58.0 1377.7 95 72.8 1407.0 100 64.5 1403.0
+50 65 85.3 100 61.4 221.4 100 56.8 1645.2 90 81.7 3356.1 95 79.0 3183.2
+60 40 90.6 100 128.2 214.2 90 71.4 2822.8 70 92.0 6337.2 95 88.7 6155.1
+70 0 - 100 65.5 2158.2 90 70.5 3158.8 60 104.5 10385.0 65 110.7 11735.3
+80 0 - 100 71.8 1405.5 80 83.8 4660.8 60 111.7 17196.1 70 108.9 20330.8
+90 0 - 95 82.8 1909.0 60 96.9 6292.6 30 117.8 29474.4 25 123.8 36040.6
+100 0 - 85 76.5 3553.8 60 97.2 6982.9 40 117.5 37877.7 50 121.0 41383.7
+128 0 - 0 - - 70 96.6 8615.5 0 128.0 - 5 127.5 129548.7
+[Uncaptioned image] random $\times64$-20 10 100 98.0 100 108.0 0.9 50 112.5 612.1 100 97.7 10.6 95 97.3 7.8
+20 90 108.8 100 115.6 6.2 10 126.5 1382.0 100 107.8 59.4 100 105.7 47.8
+30 65 112.0 100 118.2 23.4 10 126.7 2190.3 90 112.2 172.2 100 109.5 141.8
+40 60 128.5 100 122.6 45.0 0 128.0 - 100 112.8 382.4 100 112.3 328.9
+50 40 127.0 100 124.0 109.8 0 128.0 - 85 116.8 889.7 100 113.8 638.0
+60 25 134.5 100 130.2 110.8 0 128.0 - 65 120.8 1258.0 95 116.1 964.6
+70 0 - 100 128.2 183.0 0 128.0 - 50 123.5 1823.1 85 120.8 1640.2
+80 0 - 100 128.7 211.9 0 128.0 - 20 126.8 2947.4 90 120.2 2520.7
+90 0 - 95 130.4 245.4 0 128.0 - 60 124.6 4095.8 80 122.2 3187.4
+100 0 - 90 130.1 462.1 0 128.0 - 25 127.2 6134.2 45 125.3 5290.0
+128 0 - 5 132.0 607.6 0 128.0 - 0 128.0 - 15 127.7 12250.5
+[…]
+16 85 128.7 100 124.0 21.8 0 256.0 - 100 130.7 191.3 100 122.2 128.4
+32 20 148.0 100 136.7 105.2 0 256.0 - 85 163.5 1195.6 100 127.0 627.5
+64 0 - 100 153.1 759.9 0 256.0 - 90 190.7 8180.1 100 146.1 6722.3
+128 0 - 0 - - 0 256.0 - 5 254.8 112519.4 90 207.7 77108.9
+[Uncaptioned image] warehouse $1×\times6$3 8 100 175.7 100 192.8 0.8 0 256.0 - 100 191.8 4.3 100 191.6 1.7
+16 100 196.2 100 215.1 1.0 0 256.0 - 100 213.3 15.6 100 212.9 11.1
+32 15 221.3 100 220.3 10.7 0 256.0 - 100 219.2 89.0 100 218.1 57.7
+64 0 - 100 227.6 23.6 0 256.0 - 100 227.6 356.9 100 227.8 265.1
+128 0 - 100 234.7 139.1 0 256.0 - 95 236.8 2016.5 90 239.1 1502.4
+[…]
+Figure 2: Strategy usage summary on random-32-32-20 (random-32), random-64-64-20 (random-64), den312d, and warehouse maps. Values are averaged across 20 problem instances.
+[…]
 Conclusion
@@
 While our method has a central coordinator, the burden on each individual agent is drastically lower, as it does not require constant, high-bandwidth sensing of its environment.
-A.2 Literature review
+A.2 Properties of Tiered Replanning Strategies
+In this section, we summarize the soundness and (local) completeness properties of the tiered replanning operators introduced in S4 (Replanning).
+We assume a standard discrete-time MAPF model on a finite grid graph $G=(V,E)$ with a finite set of agents $\{a_{1},\dots,a_{N}\}$.
+Time is discrete, $t=0,1,\dots,T_{\max}$; each agent can either move along an edge $(u,v)\in E$ or wait at its current vertex.
+A plan for agent $a_{i}$ is a finite sequence of vertices
+$\pi_{i}=(v_{i}(0),v_{i}(1),\dots,v_{i}(T_{i})),$
+where $(v_{i}(t),v_{i}(t+1))$ is either an edge in $E$ or a wait, with $v_{i}(0)$ and $v_{i}(T_{i})$ denoting its fixed start and goal positions, respectively.
+A joint plan is the collection $\Pi=\{\pi_{1},\dots,\pi_{N}\}$.
+We adopt the standard MAPF collision model: a joint plan $\Pi$ is collision-free if it contains no vertex collision
+$\exists i\neq j,\ \exists t:\ v_{i}(t)=v_{j}(t),$
+and no edge collision
+$\exists i\neq j,\ \exists t:\ v_{i}(t)=v_{j}(t+1)\ \wedge\ v_{i}(t+1)=v_{j}(t).$
+Throughout this discussion, we assume that each tier uses an underlying search routine (single-agent planner or joint planner) that is complete on its induced finite state-time graph (e.g., BFS or A* with an admissible, consistent heuristic), and that any candidate repair is re-simulated against the other agents before being committed.
+Soundness.
+We first formalize soundness of a replanning operator.
+Definition 1 (Soundness)
+A replanning operator $\mathcal{R}$ is sound if, whenever it returns a modified joint plan $\Pi^{\prime}$, the resulting joint plan is collision-free under the given collision model.
+In our framework, S4.1–S4.4 operate under constraints provided by the control module (S3) and are only accepted if a global collision check succeeds.
+Lemma 1 (Soundness of S4.1–S4.4)
+Assume that each replanning operator (yield-based, bounded static, bounded dynamic, and local joint planning) only considers legal moves (graph edges or waits), obeys all constraints imposed by S3 (forbidden cells, reservations, and joint consistency), and that any candidate repaired plan is re-simulated against all other agents and discarded if any collision is detected.
+Then every committed repair produced by S4.1–S4.4 is collision-free; in particular, each tier is sound.
+Proof 1 (Proof sketch)
+Each operator constructs candidate paths by exploring a constrained state-time graph in which illegal moves (violating static constraints, dynamic reservations, or internal joint consistency) are excluded.
+By construction, no candidate produced by the underlying planner violates these constraints.
+Before committing any repair, the resulting joint plan $\Pi^{\prime}$ is simulated and checked for vertex and edge collisions against all other agents.
+If a collision is found, the candidate is rejected.
+Hence any repair that is actually committed has passed both constraint enforcement and a global collision check, and is therefore collision-free.
+Yield-based local coordination (S4.1).
+The yield operator chooses a single agent (possibly anchored at its start or goal) and constructs a short detour to a nearby “parking” cell, waits there while other agents traverse the conflicted region, and then rejoins the original plan.
+The search for a parking cell is restricted to a bounded local neighborhood and does not modify other agents’ plans.
+Lemma 2 (Soundness and incompleteness of S4.1)
+The yield-based operator S4.1 is sound but not complete even for resolving a single conflict: there exist MAPF instances in which a collision can be resolved by modifying agents’ plans, but no admissible yield maneuver exists within the bounded neighborhood and single-agent restriction imposed by S4.1.
+Proof 2 (Proof sketch)
+Soundness follows directly from Lemma 1: any candidate yield path is simulated and only committed if no collisions remain.
+To see incompleteness, consider a long, narrow corridor of width one, and two agents at opposite ends in the middle of the corridor, wishing to swap positions.
+A known solution exists (e.g., coordinating their timing so one waits at the opening while the other passes), but under a strict yield model that only searches for a parking cell in a small neighborhood around the collision, it is impossible to find a nearby parking position that both satisfies the local constraints and removes the conflict.
+More generally, whenever resolving a conflict requires modifying the timing or paths of multiple agents together, or using parking locations outside the bounded neighborhood, S4.1 may fail to find a repair even though a global solution exists.
+Thus, S4.1 is sound but incomplete.
+Bounded static replanning (S4.2).
+In bounded static replanning, the control module S3 selects a single agent, fixes all other agents’ plans, and designates a time window $\{t_{\mathrm{start}},t_{\mathrm{end}}\}$ along the agent’s current plan.
+A static constraint set $\Delta_{c}\subseteq V$ encodes cells that must be avoided (e.g., conflict cells), and the planner recomputes only the affected suffix of the agent’s trajectory while treating $\Delta_{c}$ as forbidden and all other agents’ plans as fixed reservations.
+This induced subproblem can be modeled as single-agent planning on a finite state-time graph whose nodes are $(v,t)$ and whose edges correspond to legal moves that respect static constraints and avoid collisions with fixed agents.
+Lemma 3 (Local completeness of S4.2)
+Consider the state-time graph induced for S4.2 over a finite horizon, with all vertices $(v,t)$ such that $v\in\Delta_{c}$ in the relevant window removed, and all nodes and edges that would collide with fixed agents pruned.
+Assume the single-agent planner used in S4.2 is complete on this graph (e.g., BFS or A* with an admissible, consistent heuristic).
+If there exists a collision-free path for the selected agent from its state at $t_{\mathrm{start}}$ to its target (goal or intermediate waypoint) within this graph, then S4.2 will find such a path and return a valid repair.
+Proof 3 (Proof sketch)
+The state-time graph is finite, as both the underlying grid and the time horizon are finite.
+Every collision-free path that respects static constraints and avoids fixed agents corresponds to a path in this pruned graph.
+A complete search algorithm on a finite graph is guaranteed to find a path to the target node if one exists.
+Therefore, under the assumed planner, S4.2 is complete for its induced subproblem.
+Bounded dynamic replanning (S4.3).
+Bounded dynamic replanning extends S4.2 by treating portions of other agents’ plans as time-varying obstacles (reservations) over a finite horizon around the conflict.
+In the corresponding state-time graph, nodes and edges that would cause vertex or edge collisions with these reservations are removed.
+Lemma 4 (Local completeness of S4.3)
+Under the same assumptions as Lemma 3, but with additional time-indexed reservations derived from other agents’ plans, bounded dynamic replanning (S4.3) is complete for its induced single-agent subproblem: if there exists a collision-free path that respects both static constraints and dynamic reservations in the finite horizon, the underlying complete planner will find it.
+Proof 4 (Proof sketch)
+The state-time graph with dynamic reservations remains finite; reservations simply prune additional nodes and edges corresponding to disallowed states and transitions.
+Any path that respects all reservations corresponds to a path in this pruned graph.
+A complete search will discover such a path if it exists.
+Hence S4.3 is locally complete for its induced single-agent subproblem.
+Local joint planning (S4.4).
+When single-agent repairs are insufficient, S4.4 jointly replans a small subset of agents $S\subseteq\{1,\dots,N\}$ over a bounded time window around the conflict, treating all other agents as fixed reservations.
+The induced local problem can be modeled as a joint state-time graph whose nodes are $(\mathbf{v}(t),t)$, where $\mathbf{v}(t)=(v_{k}(t))_{k\in S}$ is a joint configuration with no internal vertex collisions and no collisions with reserved agents.
+Lemma 5 (Local completeness of S4.4)
+Consider the joint state-time graph for a fixed subset of agents $S$ and a finite horizon $\{t_{\mathrm{start}},t_{\mathrm{end}}\}$, where nodes correspond to collision-free joint configurations of $S$ and edges correspond to legal joint moves (Cartesian products of individual moves) that do not introduce collisions among $S$ or with reserved agents.
+Assume that S4.4 runs a complete joint planner (e.g., joint A* with an admissible heuristic) on this graph.
+If there exists a collision-free joint path from the initial configuration of $S$ at $t_{\mathrm{start}}$ to the designated joint target region by $t_{\mathrm{end}}$, then S4.4 will find such a joint path.
+Proof 5 (Proof sketch)
+The joint state-time graph is finite: the number of joint configurations is bounded by $|V|^{|S|}$ per time step, and the time horizon is finite.
+Edges are finite and defined by admissible joint actions.
+A complete search algorithm on this finite graph is guaranteed to find a path to a reachable goal region.
+Any valid joint plan corresponds to such a path.
+Therefore S4.4 is complete for its induced joint subproblem.
+Discussion and overall behavior.
+The above lemmas show that each tier is sound and that the static, dynamic, and joint operators (S4.2–S4.4) are complete for their respective induced subproblems under the assumption of complete search over the chosen finite horizon.
+Yield-based coordination (S4.1) is sound but intentionally incomplete.
+In the full algorithm, these tiers are combined with a deferral mechanism: if all tiers fail for an agent within the current attempt and bounded parameters, the agent is temporarily deferred (parked and excluded from further replanning), the remaining agents are resolved, and the deferred agents are reintroduced in a second phase with relaxed limits.
+With fixed bounds on time windows, joint subset sizes, and search effort, the overall tiered strategy is best viewed as a sound, information-efficient repair framework with local completeness guarantees, rather than a globally complete MAPF algorithm.
+In principle, completeness for the full MAPF instance could be recovered by allowing the temporal windows and sizes of jointly replanned subsets to grow as needed and by repeatedly applying the tiers (including deferral) until all conflicts are resolved; however, this configuration would come at a substantially higher computational cost and is not pursued in our current implementation.
+Figure 4: Five map configuration, initial state and goal state, used for testing our IC–MAPF algorithm on TurtleBot4s’
+Figure 5: Indoor 6x6 grid setup with TurtleBot4’
+A.3 Hardware Results: TurtleBot4
+We evaluated IC–MAPF on five TurtleBot4 robots deployed in a $6\times 6$ indoor grid.
+The centralized coordinator module (S3) ran on a central host PC, while each robot also executed its own local planning (S1) and replanning procedures (S4) under isolated ROS 2 namespaces.
+Robots communicated only through a shared synchronization topic (/sync_barrier) and operated under a ROS 2 Discovery Server setup, ensuring that no robot was configured to discover or access any other robot’s topics or services; the only cross-namespace visibility permitted was the synchronization channel.
+No onboard perception (LiDAR, RGB-D, or OAK-D) was used; instead, robots followed their assigned plans using open-loop odometry control.
+All remaining ROS interfaces, including odometry and velocity commands, remained fully namespace-isolated, ensuring minimal inter-agent information exchange.
+At the start of each run, robots reset their poses via the /reset_pose service and waited for odometry stabilization.
+Execution proceeded in strict lockstep: at every step, a robot (i) rotated to the required heading using closed-loop angular control, and (ii) moved forward by $0.45$ m using odometry-based distance tracking.
+After completing its action, each robot published a synchronization message and waited until all robots reached the same step, ensuring coordinated, collision-free execution across the team.
+We evaluated five MAPF problems, each repeated five times, and logged communication statistics, execution times, and stepwise deviations from planned poses.
+Communication and execution metrics are reported in Table 4, while execution-quality statistics are summarized in Table 5.
+These experiments demonstrate that IC–MAPF can be reliably deployed on real robots with extremely lightweight sensing and communication requirements.
+In contrast, learning-based MAPF methods that depend on field-of-view observations would require additional sensors, onboard neural inference, higher energy consumption, and greater bandwidth usage.
+IC–MAPF’s ability to operate using only minimal synchronization messages highlights its practicality for resource-constrained, scalable, and privacy-preserving multi-robot systems.
+Map Plan. Msgs IU Exec. Msgs Exec. Time (s)
+Map 1 6 48 55 110.3
+Map 2 16 15 75 162.6
+Map 3 8 21 65 129.5
+Map 4 8 14 65 145.7
+Map 5 10 54 95 260.5
+Table 4: Communication and execution-related statistics for five TurtleBot4 trials.
+Map MS Pos. Err. (cm) Head. Err. ()
+Map 1 11 2.08 1.97
+Map 2 15 2.63 2.42
+Map 3 13 2.14 2.09
+Map 4 13 2.12 2.01
+Map 5 19 2.58 2.54
+Table 5: Execution quality measures: makespan, average positional deviation, and heading deviation for five TurtleBot4 runs. Harder problems (higher MS) show slightly larger systematic drift.
+A.4 Literature review
 Conflict-Based Search (CBS)
@@
 S1 – Agent Planning: Distributed: In both PRIMAL and PRIMAL2, each agent independently plans its path using policies learned through a combination of RL and IL.
-These policies are trained to enable agents to navigate towards their goals based on local observations without centralized coordination (Sartoretti et al.
-2019; Damani et al.
-2021b).
+These policies are trained to enable agents to navigate towards their goals based on local observations without centralized coordination (Sartoretti and others 2019; Damani et al.
+2021).
 S2 – Collision Detection: Distributed: Agents detect potential collisions based on their local observations of the environment.
@@
 Later Iterations: Replanning shifts to a centralized approach using PP, focusing on efficiency and resolving any remaining conflicts.
-Table 2: Related Literature Categorization into – S1: Agent Planning; S2: Collision Detection; S3: Collision avoidance policy; S4: Agent Replanning (SB: Search Based; LB: Learning Based)
+Table 6: Related Literature Categorization into – S1: Agent Planning; S2: Collision Detection; S3: Collision avoidance policy; S4: Agent Replanning (SB: Search Based; LB: Learning Based).
+Our approach is IC–MAPF with S4 variants 4.0-4.3.
 Method S1 S2 S3 S4
-Our* Decentralized Centralized Centralized Decentralized
+Our: IC–MAPF Decentralized Centralized Centralized Distributed (4.0 - 4.3)
 (SB) CBS Decentralized Centralized Centralized Centralized
@@
 (LB) LNS2+RL Centralized Centralized Hybrid Hybrid
-A.3 Training Methods
+A.5 Training Methods
 PPO Training Procedure
@@
 ResNetDQN Network Architecture for DDQN Training
-ResNetDQN is a residual‐network architecture for approximating the action‐value function $Q^{\pi}(s,a)$ in grid‐based environments trained using the DDQN algorithm (Van Hasselt, Guez, and Silver 2016).
+ResNetDQN is a residual‐network architecture for approximating the action‐value function $Q^{\pi}(s,a)$ in grid‐based environments trained using the DDQN algorithm (Van Hasselt et al.
+2016).
 It combines a deep convolutional stem with residual blocks (He et al.
@@
 Smoothed curves represent moving averages, enhancing the visibility of underlying performance trends.
-Model Comparison
-Table 3: Performance comparison of reinforcement learning methods (BFS-DQN, BFS-PPO, A*-DQN, A*-PPO) across different grid environments and agent counts.
-Results are averaged over 10 problem instances per configuration.
-Evaluated metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T), with the Time metric averaged over all trials.
-Shows that the model trained using DDQN, selected for our main paper, yields better results over PPO model.
-(a) 11x11 maze grid, $\sim$45% static obstacles (5–20 agents). Time in seconds.
-SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (s) ($\downarrow$)
-Methods 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20
-BFS-DDQN 100.00% 100.00% 98.00% 17.8 23.6 27.4 9 53 191 1.1 4.2 11.8
-BFS-PPO 100.00% 94.00% 74.00% 22.1 28.6 28.6 9 48 125 4.9 16.5 31.8
-A*-DDQN 100.00% 96.00% 82.00% 16.1 20.7 24.0 10 59 173 3.3 12.0 25.8
-A*-PPO 100.00% 96.00% 74.00% 16.6 21.4 25.2 10 60 171 3.9 14.4 30.4
-(b) 21x21 maze grid, $\sim$35% static obstacles (32, 64, and 96 agents). Time in minutes.
-SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
-Methods 32 64 96 32 64 96 32 64 96 32 64 96
-BFS-DDQN 100.00% 90.00% 0.00% 86.6 131.4 - 528 18909 - 2.7 23.9 33.5
-BFS-PPO 100.00% 0.00% 0.00% 87.6 - - 463 - - 11.2 30.2 31.1
-A*-DDQN 100.00% 80.00% 0.00% 41.4 72.0 - 457 11379 - 3.3 23.5 33.4
-A*-PPO 100.00% 0.00% 0.00% 39.6 78.3 - 407 10761 - 6.2 27.0 30.2
-(c) 25x25 warehouse grid, $\sim$24% static obstacles (32, 64, and 96 agents). Time in minutes.
-SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
-Methods 32 64 96 32 64 96 32 64 96 32 64 96
-BFS-DDQN 100.00% 100.00% 10.00% 94.3 125.9 96.0 237 6379 26553 5.2 28.0 59.5
-BFS-PPO 100.00% 40.00% 0.00% 93.6 130.0 - 231 3740 - 15.2 57.3 60.2
-A*-DDQN 100.00% 100.00% 60.00% 40.5 43.7 50.8 211 3129 16277 4.7 19.2 51.8
-A*-PPO 100.00% 100.00% 40.00% 41.6 42.4 51.5 179 3117 17382 5.7 30.0 55.4
-A.4 Additional Experiments
-We present a model comparison trained using DDQN and PPO reinforcement learning algorithms, Table 3, in Section A.3.
-This section details further empirical evaluations and an ablation study of our framework.
-An ablation study of different agent selection policies is detailed in Table 4. We have considered the information sharing setting/replanning strategy as Static & Dynamic Obstacle Replanning.
-Furthermore, a comparison of replanning settings for static and dynamic obstacles is provided in Table 5, which reflects different information sharing settings as detailed in our main paper.
-Here, we considered the agent selection policy as Fewest Future Collisions (FFC).
-We also present extended analyses on MovingAI lab’s MAPF benchmark dataset (Stern et al.
-2019b) – random map $32\times 32-10$ environment with various agent counts [8, 16, 32, 64, 96], results presented in Section A.4.
-Agent Selection Policy Comparison
-Table 4: Performance comparison of Agent Selection Policies (Random, Farthest, Fewest Future Collisions - FFC) for Alert-BFS and Alert-A* methods across different grid environments and agent counts.
-Metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T).
-Results are averaged over 10 problem instances per configuration.
-Best performance across policies for each method and agent group is bolded.
-Using Static & Dynamic obstacle replanning.
-Shows that, among the choices [random, farthest & fewest future collisions (FFC)], the FFC policy, selected for our main paper, yields best results across the different map configurations.
-(a) 11x11 maze, $\sim$45% static obstacles (5–20 agents). Time in seconds (s).
-SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (s) ($\downarrow$)
-Methods 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20
-Policy: Random Agent Selection
-Alert-BFS 98.33% 98.00% 82.00% 20.5 24.6 32.7 20 108 517 2.0 5.8 22.6
-Alert-A* 98.33% 96.00% 50.00% 17.8 23.6 24.3 16 122 259 4.3 15.7 35.6
-Policy: Farthest Agent Selection
-Alert-BFS 100.00% 94.00% 90.00% 19.8 24.6 30.3 14 91 387 1.2 6.9 17.8
-Alert-A* 98.33% 84.00% 62.00% 18.9 23.3 26.5 14 96 224 4.6 18.3 35.1
-Policy: Fewest Future Collisions (FFC)
-Alert-BFS 100.00% 100.00% 98.00% 17.8 23.6 27.4 9 53 191 1.1 4.2 11.8
-Alert-A* 100.00% 96.00% 82.00% 16.1 20.7 24.0 10 59 173 3.3 12.0 25.8
-(b) 21x21 maze grid, $\sim$35% static obstacles (32, 64, and 96 agents). Time in minutes (min).
-SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
-Methods 32 64 96 32 64 96 32 64 96 32 64 96
-Policy: Random Agent Selection
-Alert-BFS 100.00% 0.00% 0.00% 99.2 - - 1670 - - 4.1 30.1 30.4
-Alert-A* 100.00% 0.00% 0.00% 48.8 - - 931 - - 6.1 30.2 30.2
-Policy: Farthest Agent Selection
-Alert-BFS 100.00% 0.00% 0.00% 93.5 - - 1216 - - 3.1 30.2 30.2
-Alert-A* 100.00% 10.00% 0.00% 47.5 64.0 - 697 14654 - 5.3 29.8 30.3
-Policy: Fewest Future Collisions (FFC)
-Alert-BFS 100.00% 90.00% 0.00% 86.6 131.4 - 528 18909 - 2.7 23.9 30.5
-Alert-A* 100.00% 80.00% 0.00% 41.4 72.0 - 457 11379 - 3.3 23.5 30.4
-(c) 25x25 warehouse grid, $\sim$24% static obstacles (32, 64, and 96 agents). Time in minutes (min).
-SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
-Methods 32 64 96 32 64 96 32 64 96 32 64 96
-Policy: Random Agent Selection
-Alert-BFS 100.00% 20.00% 0.00% 119.0 139.0 - 722 23853 - 12.4 57.9 60.1
-Alert-A* 100.00% 100.00% 0.00% 42.4 54.6 - 302 5788 - 6.0 37.9 60.1
-Policy: Farthest Agent Selection
-Alert-BFS 100.00% 60.00% 0.00% 83.4 126.0 - 455 9836 - 8.5 47.6 60.8
-Alert-A* 100.00% 100.00% 0.00% 45.0 59.4 - 268 4644 - 6.4 34.9 60.1
-Policy: Fewest Future Collisions (FFC)
-Alert-BFS 100.00% 100.00% 10.00% 94.3 125.9 96.0 237 6379 26553 5.2 28.0 59.5
-Alert-A* 100.00% 100.00% 60.00% 40.5 43.7 50.8 211 3129 16277 4.7 19.2 51.8
-Inter-agent Information Sharing Comparison
-Table 5: Comparison of replanning settings (Only Static, Only Dynamic, Static & Dynamic Obstacles) for Alert-BFS and Alert-A* methods across different grid environments and agent counts.
-Metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T).
-Results are averaged over 10 problem instances per configuration.
-Best performance across settings for each method and agent group is bolded.
-Using FFC agent selection policy.
-Shows that, among the replanning strategies, Static + Dynamic obstacle replanning strategy (selected for our main paper), yields best results across the different map configurations.
-(a) 11x11 maze grid, $\sim$45% static obstacles (5–20 agents). Time in seconds (s).
-SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (s) ($\downarrow$)
-Methods 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20
-Setting 1: Only Static Obstacle Replanning (No inter-agent information)
-Alert-BFS 95.00% 92.00% 84.00% 17.5 23.0 25.4 8 51 205 1.4 4.6 14.2
-Alert-A* 95.00% 90.00% 78.00% 15.6 20.3 23.7 8 55 177 3.2 11.6 26.0
-Setting 2: Only Dynamic Obstacle Replanning
-Alert-BFS 63.33% 24.00% 8.00% 19.3 21.0 26.8 16 89 171 4.9 17.9 34.7
-Alert-A* 50.00% 12.00% 2.00% 14.0 13.8 15.0 6 25 49 22.9 41.4 52.1
-Setting 3: Static & Dynamic Obstacle Replanning
-Alert-BFS 100.00% 100.00% 98.00% 17.8 23.6 27.4 9 53 191 1.1 4.2 11.8
-Alert-A* 100.00% 96.00% 82.00% 16.1 20.7 24.0 10 59 173 3.3 12.0 25.8
-(b) 21x21 maze grid, $\sim$35% static obstacles (32, 64, and 96 agents). Time in minutes (min).
-SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
-Methods 32 64 96 32 64 96 32 64 96 32 64 96
-Setting 1: Only Static Obstacle Replanning (No inter-agent information)
-Alert-BFS 100.00% 10.00% 0.00% 86.6 140.0 - 528 12476 - 5.8 30.0 30.0
-Alert-A* 100.00% 20.00% 0.00% 41.4 59.0 - 457 9310 - 6.7 30.0 30.0
-Setting 2: Only Dynamic Obstacle Replanning
-Alert-BFS 0.00% 0.00% 0.00% - - - - - - 30.0 30.0 30.0
-Alert-A* 0.00% 0.00% 0.00% - - - - - - 30.0 30.0 30.0
-Setting 3: Static & Dynamic Obstacle Replanning
-Alert-BFS 100.00% 90.00% 0.00% 86.6 131.4 - 528 18909 - 2.7 23.9 30.0
-Alert-A* 100.00% 80.00% 0.00% 41.4 72.0 - 457 11379 - 3.3 23.5 30.0
-(c) 25x25 warehouse grid, $\sim$24% static obstacles (32, 64, and 96 agents). Time in minutes (min).
-SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
-Methods 32 64 96 32 64 96 32 64 96 32 64 96
-Setting 1: Only Static Obstacle Replanning (No inter-agent information)
-Alert-BFS 100.00% 100.00% 0.00% 95.8 117.4 - 230 6097 - 10.0 40.0 60.2
-Alert-A* 100.00% 100.00% 60.00% 42.2 41.6 56.0 204 3258 16694 8.9 32.3 53.3
-Setting 2: Only Dynamic Obstacle Replanning
-Alert-BFS 0.00% 0.00% 0.00% - - - - - - 61.7 62.4 60.6
-Alert-A* 0.00% 0.00% 0.00% - - - - - - 58.6 61.6 61.8
-Setting 3: Static & Dynamic Obstacle Replanning
-Alert-BFS 100.00% 100.00% 10.00% 94.3 125.9 96.0 237 6379 26553 5.2 28.0 59.5
-Alert-A* 100.00% 100.00% 60.00% 40.5 43.7 50.8 211 3129 16277 4.7 19.2 51.8
-Benchmark Dataset Evaluation
-Map Configuration: $32\times 32-10$ Random Map
-The random map $32\times 32-10$ is obtained from the Moving AI Lab’s MAPF benchmark dataset (Stern et al.
-2019b).
-Agent configurations tested include scenarios with 8, 16, 32, 64, and 96 agents.
-Each configuration was tested across 10 different problem instances (unique start and goal locations for all agents).
-The detailed performance metrics, including success rates, makespan, collisions, and computation time, are presented in Table 6.
-Table 6: Comparison of performance metrics for different methods on a $32\times 32$ random map.
-Results are averaged over 10 problem instances per configuration and presented in two subtables based on agent counts.
-Key metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T(min)), with Time averaged over all trials.
-For learning-based methods like PRIMAL and SCRIMP, the average steps taken (episode length) are presented in brackets along with average wall-clock time taken.
-For learning based methods, the environment step limit is set to 512.
-Best performing values for each metric per agent count are highlighted in bold.
-(a) 32x32 random map (8, 16, and 32 agents). The time limit per problem instance is 1 hour. For learning based methods, the environment step limit is set to 512.
-SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
-Methods 8 16 32 8 16 32 8 16 32 8 16 32
-Alert-BFS 100.00% 100.00% 100.00% 62.5 76.2 104.9 2 9 64 0.2 0.6 2.4
-Alert-A* 100.00% 100.00% 100.00% 42.7 45.5 48.1 2 8 68 0.2 0.9 2.5
-CBS 100.00% 100.00% 90.00% 40.2 43.6 47.0 1 6 35079 0.01 0.01 13.0
-ICBS 100.00% 100.00% 90.00% 40.2 43.6 47.0 1 6 15438 0.01 0.01 10.7
-PRIMAL 0.00% 0.00% 0.00% - - - - - - 0.1 (512) 0.3 (512) 1.0 (512)
-SCRIMP 100.00% 100.00% 100.00% 39.4 43.3 47.7 0 0 0 0.01 (39) 0.01 (43) 0.03 (48)
-(b) 32x32 random map (64 and 96 agents). The time limit per problem instance is 1 hour. For learning based methods, the environment step limit is set to 512.
-SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
-Methods 64 96 64 96 64 96 64 96
-Alert-BFS 100.00% 100.00% 136.5 154.4 1194 7079 12.9 33.8
-Alert-A* 100.00% 100.00% 53.2 52.0 1158 5912 13.2 30.7
-CBS 0.00% 0.00% - - - - 60.1 60.1
-ICBS 0.00% 0.00% - - - - 60.3 60.2
-PRIMAL 0.00% 0.00% - - - - 3.6 (512) 8.4 (512)
-SCRIMP 100.00% 90.00% 52.1 55.8 2 3 0.2 (52) 0.5 (101)